### v2

# LaMSUM: Creating Extractive Summaries of User Generated Content using LLMs

Abstract.
Large Language Models (LLMs) have demonstrated impressive performance across a wide range of NLP tasks, including summarization. LLMs inherently produce abstractive summaries by paraphrasing the original text, while the generation of extractive summaries – selecting specific subsets from the original text – remains largely unexplored. LLMs have a limited context window size, restricting the amount of data that can be processed at once. We tackle this challenge by introducing LaMSUM, a novel multi-level framework designed to generate extractive summaries from large collections of user-generated text using LLMs. LaMSUM integrates summarization with different voting methods to achieve robust summaries. Extensive evaluation using four popular LLMs (Llama 3, Mixtral, Gemini, GPT-4o) demonstrates that LaMSUM outperforms state-of-the-art extractive summarization methods. Overall, this work represents one of the first attempts to achieve extractive summarization by leveraging the power of LLMs, and is likely to spark further interest within the research community.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: International Conference on Web Search and Data Mining; March 10–14, 2025; Hannover, Germanyisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
Brevity is the soul of wit.
— William Shakespeare, Hamlet, Act 2, Scene 2
Social media platforms like Facebook, X (formerly Twitter), and Reddit offer a medium for individuals to express their opinions and views on various subjects, leading to a diverse array of perspectives shared through social debates, critiques, and reviews  (Zhang et al., 2024b; Chhikara et al., 2023; Dash et al., 2019a; Bražinskas et al., 2020; Zhang et al., 2016). With a vast amount of data being generated by users online, there is a growing need for summarization algorithms providing a precise and concise summary, eliminating the need for users to sift through numerous posts or reviews. Based on the resulting summary, summarization algorithms can be categorized as ‘extractive’ and ‘abstractive’. In extractive summarization, the aim is to select a subset representative of the original text (Xu et al., 2020; Liu and Lapata, 2019; Zhong et al., 2020; Zhang et al., 2022, 2023a). In contrast, abstractive summarization aims to generate natural language summaries that capture the essence of the original text, often paraphrasing the content rather than directly extracting it (Lewis et al., 2020; Zhang et al., 2020).
Extractive summarization is a crucial task with a rich body of literature, widely applied in summarizing legal cases, news articles, lectures, clinical notes, and social media content (Bhattacharya et al., 2021; Kanwal and Rizzo, 2022; Mukherjee et al., 2020; Jia et al., 2020). For example, a Google search for a topic or hashtag displays a few tweets alongside the usual lists of websites and news articles. The selection of tweets displayed in search results is similar to an extractive summarization task, where only a subset of tweets is selected (Griffin, 2015). Additionally, the Library of Congress only stores a selection of tweets as part of its archive to optimize storage space (Congress, 2017). E-commerce platforms like Amazon display only a subset of reviews in the condensed view, such that the chosen reviews are representative of all the reviews submitted by the users. Such cases are instances of extractive summarization, where the target is to select the most relevant subset that can effectively summarize a topic or a product. When summarizing user generated data, it is crucial to quote the user’s exact words, making extractive summarization particularly valuable in such cases. In the process of summarization, each post or review is treated as a basic unit – often referred to as a textual unit.
In recent years, Large Language Models (LLMs) have exhibited high performance across various tasks, including summarization (Ouyang et al., 2022; Brown et al., 2020; Tang et al., 2023b; Jin et al., 2024). Summaries generated by LLMs showcase high coherence and are overwhelmingly preferred by the human evaluators (Pu et al., 2023; Liu et al., 2024). However, using LLMs for extractive summarization has two significant limitations: (i) as generative models, LLMs naturally tend to perform abstractive summarization by paraphrasing rather than selecting the most relevant sentences; and (ii) due to the finite size of the context window, LLMs cannot handle long texts in a single input, underscoring the need for a method that allows for processing long text. Given the increasing reliance on LLMs for summarization, it is worth investigating the utility of LLMs for extractive summarization of large volumes of user-generated text.
To address these limitations, we present a novel framework LaMSUM (Large Language Model based Extractive SUMmarization) that integrates LLM-generated summaries with voting algorithms borrowed from Social Choice Theory (Brandt et al., 2016). Voting algorithms are traditionally employed in decision making processes to aggregate preferences and opinions of a large population; the prime example being the electoral system, where voters’ preferences for candidates are gathered to elect a winner that best represents the voters’ preferences. Voting systems can be categorized into single-winner and multi-winner types, depending on the number of winners. In a single-winner system, only one candidate is elected, whereas a multi-winner system allows for the selection of multiple candidates.
Extractive summarization can be viewed as a multi-winner election, where the input units act as candidates, and the units chosen for the summary are considered as the winners. A multi-winner voting system includes approval-based committee (ABC) voting, where voters approve a subset of candidates without ranking, and ranked choice voting, where voters rank candidates by preference. In LaMSUM, we utilize two approval based voting algorithms – Plurality Voting (Mudambi et al., 1996) and Proportional Voting (Lackner et al., 2023), and one ranked based voting algorithm – Borda Count (Emerson, 2013). Our judicial application of voting algorithms with a multi-level summarization framework ensures that LaMSUM outperforms the state-of-the-art fine-tuned summarization models. In summary, in this work, we make the following contributions:
  • We propose a novel framework LaMSUM which can effectively summarize large (having ¿30K tokens) collection of user generated content.
  • LaMSUM considers a multi-level summarization model that utilizes voting algorithms to combine outputs to generate robust summaries.
  • Analyse whether an ensemble model with multiple LLMs in LaMSUM can lead to improved outcomes across different voting methods.
To our knowledge, this is the first work to implement extractive summarization of large user-generated texts using LLMs by combining summarization with voting algorithms. We believe this work can spawn further research in this direction.
2. Background and Related Work
In this section, we review the relevant prior works that provide the foundation for our current research.
Figure 1. Current LLMs, by default, produce abstractive summaries. Llama3-70b-8192, despite specifically prompted for extractive summarization, generates abstractive summaries. This behavior underscores the need for a targeted approach to enable LLMs to effectively generate extractive summaries.
Text Summarization Algorithms
With the growing volume of online data, the demand for algorithms that automatically shorten and summarize texts is increasing. Automatic summarization can be approached in two ways: extractive and abstractive. In extractive summarization, a subset of input collection is selected for the summary based on their perceived quality and significance, aiming to represent a larger dataset with a concise sample. Abstractive summarization, on the other hand, generates natural language summaries that capture the most critical information from the original text. Both approaches seek to provide readers with a concise overview of the textual content. Over the years, many text summarization algorithms have been proposed in the literature; the reader can refer to (Gupta and Lehal, 2010; El-Kassas et al., 2021) for detailed surveys.
Extractive Summarization of User Generated Content
A large body of prior research has focused on summarizing long documents (Allahyari et al., 2017; Dong, 2018; Cajueiro et al., 2023). In fact, traditionally, summarization tasks have focused on summarizing a single document, such as a news article or a business report. In recent years, summarization has been increasingly applied on different types of user generated text (e.g., tweets, Facebook or Reddit posts) (Mackie et al., 2014; Inouye and Kalita, 2011; Jiang et al., 2021), where the task is to summarize short, independent posts written by many users. Several extractive summarization algorithms tailored specifically for user generated content have also been proposed (Sharifi et al., 2010; Xu et al., 2013; Dash et al., 2019b; Mukherjee et al., 2022; Rudra et al., 2018; Jones et al., 2023).
Large Language Models (LLMs) for Summarization
LLMs are now being extensively used for summarization (Brown et al., 2020; Tang et al., 2023b; Jin et al., 2024). Multiple works have proposed few-shot learning frameworks for the abstractive summarization of news, documents, webpages, and generic texts (Zhang et al., 2023b; Tang et al., 2023a; Yang et al., 2023; Bražinskas et al., 2020; Laskar et al., 2023), but their primary focus remains on short documents that can fit in the LLM context window. Researchers have also observed that human evaluators are increasingly preferring LLM-generated summaries compared to other baselines (Zhang et al., 2024a; Wu et al., 2024; Goyal et al., 2023; Zhang et al., 2023c; Liu et al., 2024). Despite the advancements, recent studies have also uncovered factual inaccuracies and inconsistencies in LLM-generated summaries (Tang et al., 2024; Tam et al., 2023; Luo et al., 2023; Laban et al., 2023).
Figure 2. LaMSUM: Multi-level framework for extractive summarization of large user-generated text. Input set $\mathcal{T}$ (level 0) is divided into $\lceil{\frac{|\mathcal{T}|}{s}}\rceil$ chunks each of size s. From each chunk a summary is produced of size $q$ (refer Figure 3), $q$ length summaries from $\lceil{\frac{|\mathcal{T}|}{s}}\rceil$ chunks are merged to form the input for the next level i.e., level 1. Iteratively the same procedure is repeated till we obtain a summary of size $k$. We set $q=k$ to ensure our algorithm can effectively handle the worst-case scenario where all the textual units in the final summary may come from the same input chunk (Section 3.2).
Extractive Summarization through LLMs: The Current State By default, LLMs produce abstractive summaries, meaning that the summary text is distinct from the input text, even when it is instructed to do otherwise. To illustrate this, we present a small example in Figure 1. An LLM, when prompted, could clearly explain extractive summarization, yet, when we instructed it to perform extractive summarization on a set of 100 sentences, it fails to do so and instead generated an abstractive summary. Prior to our current work, only two studies attempted to perform similar tasks. Zhang et al. (2023b) attempted summarization of short news articles using GPT 3.5, while Chang et al. (2024) attempted abstractive summarization for book-length documents. However, both these approaches suffer from practical limitations such as lack of contextual dependencies in user generated text and the problem with positional bias in LLMs.
To the best of our knowledge, ours is the first attempt to perform extractive summarization on a large collection of user generated texts through LLMs, while tackling the challenge of positional bias. We describe our proposal in detail in the next section.
3. Generating Extractive Summaries through LLMs
In this section, we define the problem statement formally and introduce our novel summarization framework LaMSUM (Large Language Model based Extractive SUMmarization) that leverages LLMs to summarize large user-generated text.
3.1. Task Formulation
[…]
3.2. Multi-Level Summarization
LLMs have a limited context window, making it impossible to input large text collections all at once. While recent models like GPT-4 support context windows of up to 128k tokens, they still cannot accommodate book-length inputs within a single window. Consequently, the input must be divided into smaller chunks to perform the desired task (Chang et al., 2024). Thus, LaMSUM employs a multi-level framework for extractive summarization, enabling it to consider input data of any size (detailed in Figure 2)  [^1] Code is available at https://anonymous.4open.science/r/LaMSUM/ .
The set $\mathcal{T}$, which contains the original textual units, is provided as input at level 0 and is divided into $\lceil{\frac{|\mathcal{T}|}{s}}\rceil$ number of chunks of size $s$. From each chunk of size $s$, we generate a summary (discussed in Section 3.3) of size $q$ (where $q<s$), and repeat this process for all $\lceil{\frac{|\mathcal{T}|}{s}}\rceil$ chunks.  [^2] Note that a chunk of size $s$ refers to a chunk containing $s$ textual units. Likewise, a summary of size $q$ indicates a summary of $q$ textual units. $|\mathcal{T}|$ denotes the number of textual units present in $\mathcal{T}$. We then merge all these $q$ length summaries obtained from level 0 to form an input for the next level i.e., level 1. We repeatedly perform this process until we obtain the final summary of length $k$. Note that the last chunk maybe less than $q$ in size, in such case we move all the textual units of the respective chunk to the next level (refer Algorithm 1).
An alternate strategy would be to divide the input $\mathcal{T}$ into $\frac{|\mathcal{T}|}{s}$ chunks each of size $s$ and from each chunk select $\frac{k\cdot s}{|\mathcal{T}|}$ sentences to be included in the summary. However, this approach assumes a uniform distribution of potential candidates across chunks that can be included in the final summary. In LaMSUM, we keep $q=k$ i.e., we extract $k$ textual units from each chunk eliminating the chance of missing any potential candidate (discussed in Section 4.4). In the worst-case scenario, all $k$ units in the final summary can come from a single chunk, and our algorithm can handle such cases effectively.
It is important to note that we are dealing with user-generated content, such as tweets, which lack contextual connections. Unlike book summarization, where chapters are interconnected and the context of previous chapters is crucial for summarizing the current one, tweets are generally standalone and contextually independent. Thus, our approach of independently deriving summaries from each chunk works well in our setup, as each textual unit operates independently of the others and there are no long-range dependencies.
Figure 3. Textual units (e.g., posts) in the input chunk are shuffled to account for the positional bias. $m$ different chunk variations are obtained through shuffling, which are subsequently summarized using LLMs. $m$ summaries are then aggregated by voting algorithms to get the final summary.
3.3. Summarizing a Chunk
Next, we discuss how LaMSUM summarizes a chunk (Algorithm 2) by tackling the positional bias in LLMs and leveraging voting algorithms drawn from Social Choice Theory (Brandt et al., 2016).
3.3.1. Tackling Positional Bias
Prior research  (Brown and Shokri, 2023; Zhang et al., 2023b; Jung et al., 2019; Wu et al., 2024) has highlighted that summarization using LLM is prone to positional bias, i.e., the sentences located in certain positions, such as the beginning of articles, are more likely to be considered in the summary. To address this issue and generate a robust summary, we create $m$ different variations by shuffling the textual units within the input chunk. This ensures that each unit has the opportunity to appear in different positions within the input text (refer Figure 3).
3.3.2. Zero-shot prompting
For each input chunk, we obtain $m$ different summaries (one for each variation) by prompting the LLM. We employ the following two prompts to obtain the summaries –
a
Select the most suitable units that summarize the input text.
Prompt: ”Input consists of ¡chunk_size¿ sentences. Each sentence is present in a new line. Each sentence contains a sentence number followed by text. You are an assistant that selects best ¡summary_length¿ sentences (subset) which summarizes the input. Think step by step and follow the instructions. ¡sentences¿”
b
Generate a ranked list in descending order of preference.
Prompt: ”Input consists of ¡chunk_size¿ sentences. Each sentence is present in a new line. Each sentence contains a sentence number followed by text. You are an assistant that outputs the sentences in the decreasing order of their relevance to be included in the summary. Think step by step and follow the instructions. ¡sentences¿ Remember that output should contain all the sentences in the decreasing order of their preference.”
Algorithm 1 Algorithm for multi-level summarization
Input: ${\mathcal{T},k,s,q,m}$
${}SS{\SS=\{\}}$triangleright $SS\SS$ stores the final summary
[…]
     $n_{chunks}=\lceil{\frac{|\mathcal{T}|}{s}}\rceil$triangleright number of chunks in set $\mathcal{T}$
     $L=\{\}$triangleright $L$ stores the results of a given level
     for $i\leftarrow 1$ to $n_{chunks}$ do
         $si=(i-1)*s$triangleright starting index of chunk
         $ei=i*s$triangleright ending index of the chunk
         if $i=n_{chunks}$ then $\triangleright$ if last chunk
              $ei=|\mathcal{T}|$triangleright ending index is equal to length of $\mathcal{T}$
         end if
         $width=ei-si$triangleright number of textual units in a chunk
         if $width<=q$ then $\triangleright$ if last chunk
              $L=L\cup t_{si}\cup t_{si+1}\cup...\cup t_{ei-1}$
$\triangleright$ add all textual units to the result
         else
              $L=L\cup\textsc{ChunkResult}(\mathcal{T},si,ei,q,m)$
$\triangleright$ add summary of each chunk to result $L$
         end if
     end for
     $\mathcal{T}=L$triangleright update the input $\mathcal{T}$ for the next level
[…]
end while
Output: $SS\SS$
Algorithm 2 Algorithm for summarization of a chunk
function ChunkResult($\mathcal{T},si,ei,q,m$)
     $X=\{\}$
     for $i\leftarrow 1$ to $m$ do $\triangleright$ for each variation of a chunk
         $V=\textsc{Shuffle}(\mathcal{T},si,ei,i)$triangleright shuffle with random state $i$
         $R=\textsc{LLM}(V,q)$triangleright obtain summary from the LLM
         $C=\textsc{Check}(R,\mathcal{T},si,ei)$triangleright output calibration
         $X.add(C)$
     end for
     return $\textsc{VOTING}(X,q)$triangleright voting to obtain the final summary
end function
 Original Post LLM Modified Output
  My allergies be acting up &amp; Claritin be letting them &gt;,&lt; My allergies be acting up & Claritin be letting them.
I spent the extra money to have fast relief, but instead I am SUFFERING #gimmemymoneyback I spent the extra money to have fast relief, but instead I am SUFFER
Table 1. Examples illustrating that LLMs when selecting textual units for summarization, often demonstrate a propensity to alter certain words or introduce new ones.
3.3.3. Output Calibration
LLMs may alter certain words from the input text while generating extractive summaries, as shown in Table 1. Thus, we perform additional checks to ensure that the textual units selected in the summary are indeed a subset of $\mathcal{T}$. If the post selected by the LLM (say $x$) is not present in the original text $\mathcal{T}$, we identify the post with the closest resemblance to $x$ by computing the edit distance (Ristad and Yianilos, 1998). LLMs may also hallucinate, generating new sentences rather than selecting units from the input. In such instances, the edit distance between the generated unit $x$ and all the original textual units would be high. We discard these sentences from the output.
3.4. Reimagining Summarization as an Election
As mentioned earlier, for a given chunk, we obtain $m$ summaries – one for each variation. We imagine the process of creating the final summary from these $m$ summaries to be a multi-winner election, where the textual units in $m$ summaries correspond to ballots (candidates) and the role of the voting algorithm is to pick $q$ winners. We employ three different voting methods, namely Plurality Voting (Mudambi et al., 1996), Proportional Approval Voting (PAV) (Lackner et al., 2023) and Ranked Choice Voting (Emerson, 2013) to determine the final summary. Due to the varying input requirements of different voting methods, changing both the prompting approach and the output generated by the LLM becomes imperative.
Plurality voting and proportional voting are approval-based voting methods where voters can select multiple candidates they approve of without indicating a specific preference order. In multi-winner plurality voting (also known as block voting), each voter casts multiple votes and the candidates are selected based on the number of votes polled. In the context of summarization, a textual unit is treated as a candidate, and the LLM acts as the voter. We select the textual units in the decreasing order of the votes polled, till we obtain a summary of size $q$. PAV evaluates the satisfaction of each voter in the election outcome. A voter’s satisfaction is measured based on – amongst the number of candidates they voted for, how many are selected in the election. In the realm of summarization, PAV selects the textual units based on the amount of support each unit receives in $m$ summaries. Since both plurality and proportional are approval-based voting algorithms, the units are either approved or disapproved by the underlying LLM, with no explicit ranking or preference order. In this case, we prompt the LLM to select the best ¡$q$¿ sentences that summarize the input text as shown in Section 3.3.2 a.
On the other hand, ranked choice voting entails assigning a score to each textual unit and subsequently selecting the highest-scoring units for inclusion in the summary. For ranked voting, we use the Borda count, a positional voting algorithm (Emerson, 2013). In the Borda method, each candidate is assigned points corresponding to the number of candidates ranked below them: the lowest-ranked candidate receives 0 points, the next lowest gets 1 point, and so forth. The candidates with the highest aggregate points are declared as the winners. In ranked voting, we prompt the LLM to output sentences in descending order of their suitability for the summary as discussed in Section 3.3.2 b.
It is important to note that the prompting technique and the output generated by LLM vary for different voting methods. In approval voting the output from LLM is a list of $q$ textual units that LLM finds best suited to be included in the summary. Whereas in ranked choice voting, the output from LLM is a list of the same length as input i.e. $s$ with all the units sorted in decreasing order of their preference towards the summary, and Borda Count (Emerson, 2013) is used to identify the top $q$ textual units. In the next section, we highlight how the voting-based summarization schemes outperform the Vanilla setup, which does not use voting.
4. Experimental Setup
      Parameters Claritin US-Election MeToo
      #TextualUnits ($|\mathcal{T}|$) 3998 2107 483
#InputWords 53609 35522 16737
AV #CTU ($s$) 200 150 75
#CSTU ($q$) 100 100 50
RV #CTU ($s$) 40 40 40
#CSTU ($q$) 20 20 20
#SummaryTU ($k$) 100 100 50
Table 2. Input parameters used for the proposed framework LaMSUM. #TextualUnits is the number of textual units in the input set i.e. $|\mathcal{T}|$. #InputWords represents the number of words present in the input set $\mathcal{T}$. #CTU is the number of textual units in a chunk i.e. chunk size represented as $s$. #CSTU is the number of textual units in the chunk summary represented as $q$. #SummaryTU is the number of textual units present in the final summary i.e. $k$. AV and RV represents Approval Voting and Ranked Voting respectively.
4.1. Dataset
Our experiments are conducted on three publicly available datasets, consisting of crowd-sourced data from X, listed in Table 2 (Dash et al., 2019a). The Claritin dataset contains 3,998 tweets about the benefits and the side-effects of the anti-allergic drug Claritin. US-Election dataset contains 2,107 tweets from 2016 US Presidential Election where people support and attack different political parties. Me-Too dataset includes 483 tweets from the October 2018 MeToo movement, where individuals recount the harassment cases they experienced. [^3] Dataset: https://github.com/ad93/FairSumm/tree/master/Dataset During pre-processing, we remove all web links and duplicate entries from the datasets. The datasets also include gold-standard summaries (reference summaries) which are human-generated i.e., the textual units that are strong candidates for inclusion in the summary are selected by the humans. Claritin and US-Election dataset, each have three gold standard summaries, comprising 100 textual units each. Me-Too dataset has two gold-standard summaries with 50 textual units each.
4.2. Large Language Models (LLMs)
LLMs are characterized by their extensive parameter sizes and remarkable learning abilities (Zhao et al., 2023; Chang et al., 2023). In our work, we utilize three open-source LLMs and one proprietary LLM to conduct experiments: llama3-70b-8192 from Meta (Touvron et al., 2023), mixtral-8x7b-32768 from Mistral AI (Jiang et al., 2024), gemini-1.0-pro from Google (Gemini Team, 2023) and
gpt-4o-mini-2024-07-18 from OpenAI (OpenAI, 2024). Across all experiments, we keep temperature, top probability and output tokens as 0, 0.8 and 8192 respectively.
4.3. Evaluation Metric
For evaluating the quality of summaries generated by LaMSUM, we report ROUGE-1, ROUGE-2, and ROUGE-Lsum scores (Lin, 2004). ROUGE-1, ROUGE-2 and ROUGE-L respectively evaluate the overlap of unigrams, bigrams and longest common subsequence between the generated summary and the reference summary. ROUGE-Lsum is more suitable for extractive summarization, as it applies ROUGE-L at sentence level and then aggregates all the results to obtain the final score.
4.4. LaMSUM Input Parameters
Input parameters for LaMSUM (Algorithm 1), such as $|\mathcal{T}|$ (total number of textual units in the set), $s$ (chunk size) and $k$ (length of summary) are listed in Table 2 for different voting algorithms and datasets. The value of $m$ (number of shuffling) for all the datasets was set to 5.
If $q\in[k,s)$, our proposed method can handle worst case scenario where all the textual units in the final summary may originate from a single chunk of level 0. As $q$ approaches $s$, more levels are required to converge to the final summary. The optimal value of $q$ that can handle worst case and also reduce the number of levels in multi-level summarization is $k$, thus we keep $q=k$ for experiments with approval voting. For instance, if $s$ is 200 and $q$ is 100, this indicates that only 50% of the units from each chunk advance to the next level.
In ranked voting algorithm, we maintain smaller value for chunk size ($s$) to ensure that the LLMs output, which is of size $s$, fits within the context window. Additionally, as chunk size increases, LLM often does not output all the sentences, instead produce generalized statements like “similarly for other sentences we find the rank”. Therefore, it is essential to keep the chunk size smaller. For ranked voting, we set $s$ and $q$ to 40 and 20 respectively, upholding the selection ratio of 50% at each level.
5. Experimental Evaluation
In this section, we present the empirical comparison of LaMSUM with competent baseline models and voting algorithms across datasets (refer Table 3).
                      Models Claritin US-Election         MeToo
R1 R2 RLSum R1 R2 RLSum R1 R2 RLSum
                     ClusterRank 50.08 14.17 48.64 55.18 13.09 53.51 55.81 55.81 52.48
DSDR 28.01 8.40 27.45 48.04 9.53 46.44 55.03 17.30 52.04
LexRank 45.04 19.71 44.74 42.63 10.78 41.64 42.70 11.32 40.91
SummBasic 58.25 19.29 56.76 55.36 12.43 53.94 58.23 21.53 56.07
GPT2 61.61 23.58 60.74 55.86 15.07 54.81 40.63 11.24 38.86
BERT 57.30 22.37 56.21 55.89 15.44 55.00 45.72 10.76 43.50
XLNET 55.52 21.37 54.75 56.48 15.72 55.41 36.58 08.50 34.48
BERTSUM (Liu and Lapata, 2019) 57.87 22.75 55.96 59.00 17.51 57.41 57.11 23.08 54.84
 Vanilla Llama3 58.07 21.86 56.73 56.12 12.93 54.71 51.84 17.84 49.47
Mixtral 57.80 20.14 55.81 57.11 14.20 55.79 49.63 11.29 46.16
Gemini 49.70 18.33 48.78 51.00 12.36 49.19 45.51 8.95 42.51
GPT-4o mini 62.93 24.82 61.08 55.06 15.01 53.88 36.30 7.05 34.54
  Plurality Voting Llama3 61.28 23.79 59.54 60.11 18.26 58.99 55.51 18.79 52.85
Mixtral 59.13 22.90 57.40 59.55 16.01 58.12 55.41 17.62 52.63
Gemini 55.43 19.32 53.35 58.83 16.25 57.54 57.95 21.47 54.87
GPT-4o mini 64.20 26.71 62.66 58.15 15.91 56.78 54.55 19.66 51.9
Proportional Voting Llama3 61.31 23.36 59.73 58.90 15.99 57.55 58.14 14.99 55.46
Mixtral 60.30 24.03 58.58 58.46 15.74 57.28 57.36 22.97 54.82
Gemini 61.77 19.83 59.81 57.92 15.85 56.49 54.02 13.95 50.90
GPT-4o mini 64.13 26.75 62.30 58.68 19.08 57.62 48.55 13.23 45.35
 Borda Count Llama3 51.87 15.93 49.61 51.95 13.49 50.60 53.35 16.99 50.38
Mixtral 56.07 22.28 54.65 48.71 13.37 47.48 53.24 18.80 50.96
Gemini 51.45 16.92 49.39 54.01 13.38 52.56 50.13 21.45 47.96
GPT-4o mini 58.01 18.08 56.05 53.10 15.94 51.63 50.10 17.82 47.75
Table 3. Table showing metric scores from different models for various datasets. Here, R1 = ROUGE-1 Score, R2 = ROUGE-2 Score, RLSum = ROUGE-LSum Score. The best value per dataset is shown in bold and clearly Approval voting outperforms all the other methods across all the evaluation measures.
5.1. Baseline Comparison
We compare LaMSUM with the pre-neural models (ClusterRank (Garg et al., 2009), DSDR (He et al., 2012), LexRank (Erkan and Radev, 2004), SummBasic (Nenkova and Vanderwende, 2005)), transformer based models (GPT2 (Radford et al., 2019), BERT (Miller, 2019), XLNET (Yang et al., 2019)) and with fine-tuned BERTSUM (Liu and Lapata, 2019) model. As presented in Table 3 it is observed that LaMSUM outperform the state-of-the-art summarization models. Earlier work (Zhang et al., 2023b) reported that the ChatGPT model achieves lower ROUGE scores on CNN/DM and XSum dataset. But our results demonstrate that our proposed framework LaMSUM performs significantly better than other fine-tuned models for large user-generated text.
Results from the LLM exhibit variability when executed multiple times over the same input. We conducted experiments with LLMs for five iterations for the same input. Table 3 contains the maximum ROUGE score obtained from these five iterations and Table 6 (in supplementary material) displays the variance in the ROUGE scores when executed for five times.
Algorithm 3 Algorithm for summarization of a chunk in Vanilla LLM
function ChunkResult($\mathcal{T},si,ei,q,m$)
     $R=\textsc{LLM}(\mathcal{T},si,ei,q)$triangleright $q$ textual units from $\mathcal{T}\in[si,ei]$
     $C=\textsc{Check}(R,\mathcal{T},si,ei)$triangleright output calibration
     return $C$
end function
5.2. Vanilla LLM vs. LaMSUM
Our proposed framework, LaMSUM, ensures robust summary generation by shuffling and employing a voting algorithm to select the best textual units for the summary. It is crucial to compare LaMSUM with a multi-level LLM that does not use shuffling and voting, which we call Vanilla LLM. Algorithm 3 outlines the steps used by vanilla LLM to find the chunk summary. Table 3 demonstrates that the vanilla multi-level LLM has lower ROUGE scores for each LLM compared to the proposed framework LaMSUM, indicating that shuffling and voting enhances the performance.
5.3. Which Voting Algorithm Holds the Lead?
We experimented with three voting algorithms, two approval-based and one ranked-based. Experimental results indicate that LLMs with approval voting perform the best compared to the ranked voting algorithm. We hypothesized that rank-based voting would yield better results, as it makes more informed decisions about the potential sentences to be included in the summary. Contrary to our expectations, rank-based algorithms performed even worse than neural and transformer-based models. This can be attributed to multiple factors: (i) LLMs hallucinate and output sentences in the same or in the reverse order as they were in the input. (ii) Occasionally, LLMs do not output all the sentences from the input, resulting in the padding of left-out sentences towards the end of the list, which disturbs the ranking and potentially affects the result. To overcome these problems, we kept the chunk size low as discussed in Section 4.4, but the results still did not surpass those of the approval-based voting algorithm.
Takeaway: LLMs, when prompted to select sentences that can summarize the input, perform better than when tasked to rank the sentences in the order of their preference towards the summary.
5.4. What Fails to Deliver Results?
To ensure extractive summarization, we tested an additional approach – each sentence is tagged with a sentence number, LLM is prompted to select the best $q$ sentences and output only the sentence numbers of the best $q$ sentences. Thereafter, the sentences corresponding to the sentence numbers can be retrieved. For instance, if $s$ is 200 and $q$ is 100, the task is to output the sentence numbers of the best 100 sentences from a pool of 200 sentences. In such cases, LLMs hallucinate and provide an output consisting of either all the odd number sentences or all the even number sentences.
Takeaway: For extractive summarization, relying solely on indexes may result in hallucination, underscoring the importance of emitting the input content and not the numbers.
6. Ensembling of LLMs in LaMSUM
Figure 4. Ensembling applied at chunk using all the four models. $q$ length summaries are obtained from each chunk through different LLMs using the shuffling method discussed in Section 3.3. All the summaries are ensembled through voting algorithms to achieve the final summary of length $q$. This final $q$ length summary is thereafter forwarded to the next level.
 Ensemble Voting Method R1 R2 RLSum
 All 4 models Plurality 54.16 20.79 53.00
Proportional 54.70 19.95 53.14
Ranked 51.78 16.97 50.24
Best 3 models Plurality 58.88 21.74 57.60
Proportional 58.04 22.78 56.71
Ranked 57.22 19.51 55.33
Weak 3 models Plurality 50.64 18.82 49.31
Proportional 52.92 19.23 51.07
Ranked 47.50 16.74 46.37
Table 4. Results obtained through ensembling of LLMs in LaMSUM for Claritin dataset. We compare 3 different cases: i) ensemble of all 4 models – Llama, Mixtral, Gemini and GPT-4o mini ii). ensemble of best 3 models – Llama, Mixtral and GPT-4o mini iii). ensemble of weak 3 models – Llama, Mixtral and Gemini.
Different LLMs can be considered as distinct experts, each offering their vote or opinion about the textual unit to be included in the summary. In LaMSUM, each expert (LLM) operated independently, determining which sentences were worthy of inclusion. We hypothesize that using an ensemble of LLMs in LaMSUM, where these experts collaborate and reach a consensus, has potential to improve performance of the overall system (Verga et al., 2024; Schoenegger et al., 2024).
We apply ensembling at each chunk as shown in Figure 4. For a given chunk $i$ containing $s$ textual units, the goal is to produce a final summary of size $q$ by gathering opinions of all the experts. Ensembling takes place in two stages. Stage 1 - For each chunk we obtain $q$ length summary through different LLMs by using shuffling and voting method as demonstrated in Section 3.3. Stage 2 - To ensemble the $q$ length summaries from different LLMs, we again leverage the voting methods. In stage 2 we make use of the same voting method that was used in stage 1 to maintain consistency. The summary generated from the ensembling process is treated as the final summary for chunk $i$ and is carried forward to the next level.
As shown in Table 3, for Claritin dataset, GPT-4o mini outperforms the other models, while Gemini delivers weaker results. We evaluate three different scenarios based on the number of LLMs used in the ensembling process: i) all four models – Llama, Mixtral, Gemini and GPT-4o mini, ii) the top three models – Llama, Mixtral and GPT-4o mini, iii) the three weaker models – Llama, Mixtral and Gemini. Table 4 shows the results of ensembling in LaMSUM for different scenarios and voting methods.
In the approval-based voting algorithm, using all four models or just the top three models for ensembling doesn’t significantly improve performance; the resulting performance after ensembling remains close to that of the weakest model in each case. However, ensembling with the ranked-based voting algorithm yielded interesting results. When ranked voting is applied to individual LLMs, it delivers the worst performance compared to other voting algorithms, as shown in Table 3. When using ranked voting with all four models in the ensembled LaMSUM, the performance is close to that of the weakest model – Gemini (49.39) ¡ Llama (49.61) ¡ Ensemble (50.24) ¡ Mixtral (54.65) ¡ GPT-4o mini (56.05). However, when considering the ensembled LaMSUM with the top three models – Gemini (49.39) ¡ Llama (49.61) ¡ Mixtral (54.65) ¡ Ensemble (55.33) ¡ GPT-4o mini(56.05), ensemble outperforms Mixtral, and the ROUGE score increases from 50.24 to 55.33. While the ensembled LaMSUM with the top three models does not surpass GPT-4o mini, it does enhance the performance of the overall setup.
Takeaway: Ensemble of LLMs in LaMSUM yield good results with rank-based voting method, and can help improve the performance of the overall setup if we select models for ensembling that perform well individually in LaMSUM. Including all models in the ensemble could result in the weakest model becoming the bottleneck.
7. Does Fine-Tuning Help?
The performance of LLMs on specific downstream tasks, such as extractive summarization, can be suboptimal when not explicitly fine-tuned. LLMs tend to generate erroneous information while sounding persuasive and assured (Bang et al., 2023; Xu et al., 2024). In Section 2, we highlighted instances where LLMs exhibit nonplussed responses by generating abstractive summaries despite being prompted for extractive summarization. They also alter some words, so we perform output calibration in LaMSUM. To address these challenges, we explore the efficacy of instruction fine-tuning through explicit instructions (Wei et al., 2022) and compare it with LaMSUM.
We utilize the Claritin dataset for instruction fine-tuning as it has the highest number of tweets among all the three datasets. It is divided into training (60%), validation (10%), and test (30%) set. The US-Election and MeToo datasets are retained as additional test sets for evaluation. Instruction fine-tuning involves three key components: i) Instruction: The prompts used for zero-shot prompting, as detailed in Section 3.3.2. ii) Input: Set of input text $\mathcal{T}$ to be summarized. iii) Output: The reference summary comprises sentences extracted from $\mathcal{T}$.
We fine-tune an open-source LLM, llama3.1-8b from Meta (Dubey and et al., 2024), using a 15 GB Tesla T4 GPU. During the fine-tuning phase, we set the learning rate, warmup ratio, and batch size to 2e-4, 0.0, and 2, respectively. For this process, we employ PEFT (Parameter Efficient Fine Tuning), which involves freezing the layers of the pre-trained model and only fine-tuning the last few layers specific to the downstream task. We follow the QLoRA (Dettmers et al., 2023) method in PEFT training. QLoRA is an optimized variant of LoRA (Hu et al., 2021) that reduces the precision of weight parameters to 4-bit. This reduction in precision decreases the model size, which is advantageous in scenarios with limited memory available for fine-tuning.
Table 5 shows the results of the fine-tuned llama3.1-8b model. We observed that the outputs are robust and adhere closely to the instructions provided. The generated summaries contain the exact sentences extracted from the input text, but the performance of the fine-tuned model is inferior compared to LaMSUM.
 Dataset R1 R2 RLSum
 Claritin 18.60 3.11 17.14
US-Election 26.55 14.81 25.05
MeToo 28.94 10.81 24.80
Table 5. Metric scores from fine-tuned llama3.1-8b for various datasets.
Takeaway: Fine-tuning an LLM with the appropriate instructions, input, and output can facilitate extractive summarization. But fine-tuning requires a substantial amount of data with gold-standard summaries. With less data, the model may overfit, resulting in poor performance, as shown in the Table 5. Additionally, a model fine-tuned on one dataset may not perform well on another dataset. Our proposed LaMSUM addresses these issues by being effective when the input data is large enough to exceed a single context window but small enough to be used for fine-tuning.
8. Concluding Discussion
This work marks an early attempt to achieve extractive summarization of large user-generated text that exceeds a single context window using zero-shot learning. The proposed multi-level framework LaMSUM leverages approval based and ranked based voting algorithms to generate robust summaries. Experiments conducted on three distinct crowd-sourced datasets demonstrated the efficacy of LaMSUM, as it outperformed the results achieved by state-of-the-art fine-tuned models. Ensembling of LLMs in our proposed framework can enhance the performance of weaker models.
Note that there can be a concern regarding the potential data leakage, as the experiments involve newer LLMs that may have been exposed to the experimented datasets during their pre-training phase. In Section 5.2 we showcased that the vanilla LLM, which also includes LLM underperformed, whereas our proposed framework which generates robust summaries yielded good results. This highlights the efficacy of our model, even when it is exposed to data leakage.
Limitation
Our proposed framework, LaMSUM, very well handles text of any length, conditioned on the fact that the final summary fits within a single context window. Some modifications to LaMSUM may be necessary when the output summary exceeds the size of a single context window. We focused only on the user-generated text where each textual unit is independent of the others; future research could extend the framework to summarization tasks involving contextual dependency such as book summarization.
Ethical Considerations
Our research focuses on using LLMs to produce extractive summaries for user-generated text. Given the recent rise of LLMs and the growing interest in applying them across different research fields, we believe this research direction can help unveil the potential of LLMs. Below, we outline concerns that need to be considered and addressed in this research area:
  • Bias and discrimination: LLMs are biased towards their training data, which can lead them to favor certain textual units when creating summaries.
  • Transparency and accountability: LLMs are black boxes, with an opaque decision-making process, making it difficult to discern how or why specific textual units are chosen for summarization. This lack of transparency can create ethical issues, especially in scenarios where clear explanations for decisions are needed.
  • Environmental impact: The training and operation of LLMs require substantial computational resources, which contribute to increased energy consumption and may have a negative environmental impact.
While LLMs can produce high-quality summaries, their use must be approached with careful consideration of potential ethical implications.

### v4

@@
-LaMSUM: Creating Extractive Summaries of User Generated Content using LLMs
-Abstract.
-Large Language Models (LLMs) have demonstrated impressive performance across a wide range of NLP tasks, including summarization.
-LLMs inherently produce abstractive summaries by paraphrasing the original text, while the generation of extractive summaries – selecting specific subsets from the original text – remains largely unexplored.
-LLMs have a limited context window size, restricting the amount of data that can be processed at once.
-We tackle this challenge by introducing LaMSUM, a novel multi-level framework designed to generate extractive summaries from large collections of user-generated text using LLMs.
+LaMSUM: Amplifying Voices Against Harassment through LLM Guided Extractive Summarization of User Incident Reports
+Abstract
+Citizen reporting platforms like Safe City in India help the public and authorities stay informed about sexual harassment incidents.
+However, the high volume of data shared on these platforms makes reviewing each individual case challenging.
+Therefore, a summarization algorithm capable of processing and understanding various Indian code-mixed languages is essential.
+In recent years, Large Language Models (LLMs) have shown exceptional performance in NLP tasks, including summarization.
+LLMs inherently produce abstractive summaries by paraphrasing the original text, while the generation of extractive summaries – selecting specific subsets from the original text – through LLMs remains largely unexplored.
+Moreover, LLMs have a limited context window size, restricting the amount of data that can be processed at once.
+We tackle these challenge by introducing LaMSUM, a novel multi-level framework designed to generate extractive summaries for large collections of Safe City posts using LLMs.
 LaMSUM integrates summarization with different voting methods to achieve robust summaries.
-Extensive evaluation using four popular LLMs (Llama 3, Mixtral, Gemini, GPT-4o) demonstrates that LaMSUM outperforms state-of-the-art extractive summarization methods.
-Overall, this work represents one of the first attempts to achieve extractive summarization by leveraging the power of LLMs, and is likely to spark further interest within the research community.
-copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: International Conference on Web Search and Data Mining; March 10–14, 2025; Hannover, Germanyisbn: 978-1-4503-XXXX-X/18/06
-1. Introduction
-Brevity is the soul of wit.
-— William Shakespeare, Hamlet, Act 2, Scene 2
-Social media platforms like Facebook, X (formerly Twitter), and Reddit offer a medium for individuals to express their opinions and views on various subjects, leading to a diverse array of perspectives shared through social debates, critiques, and reviews  (Zhang et al., 2024b; Chhikara et al., 2023; Dash et al., 2019a; Bražinskas et al., 2020; Zhang et al., 2016).
-With a vast amount of data being generated by users online, there is a growing need for summarization algorithms providing a precise and concise summary, eliminating the need for users to sift through numerous posts or reviews.
-Based on the resulting summary, summarization algorithms can be categorized as ‘extractive’ and ‘abstractive’.
-In extractive summarization, the aim is to select a subset representative of the original text (Xu et al., 2020; Liu and Lapata, 2019; Zhong et al., 2020; Zhang et al., 2022, 2023a).
-In contrast, abstractive summarization aims to generate natural language summaries that capture the essence of the original text, often paraphrasing the content rather than directly extracting it (Lewis et al., 2020; Zhang et al., 2020).
-Extractive summarization is a crucial task with a rich body of literature, widely applied in summarizing legal cases, news articles, lectures, clinical notes, and social media content (Bhattacharya et al., 2021; Kanwal and Rizzo, 2022; Mukherjee et al., 2020; Jia et al., 2020).
-For example, a Google search for a topic or hashtag displays a few tweets alongside the usual lists of websites and news articles.
-The selection of tweets displayed in search results is similar to an extractive summarization task, where only a subset of tweets is selected (Griffin, 2015).
-Additionally, the Library of Congress only stores a selection of tweets as part of its archive to optimize storage space (Congress, 2017).
-E-commerce platforms like Amazon display only a subset of reviews in the condensed view, such that the chosen reviews are representative of all the reviews submitted by the users.
-Such cases are instances of extractive summarization, where the target is to select the most relevant subset that can effectively summarize a topic or a product.
-When summarizing user generated data, it is crucial to quote the user’s exact words, making extractive summarization particularly valuable in such cases.
-In the process of summarization, each post or review is treated as a basic unit – often referred to as a textual unit.
-In recent years, Large Language Models (LLMs) have exhibited high performance across various tasks, including summarization (Ouyang et al., 2022; Brown et al., 2020; Tang et al., 2023b; Jin et al., 2024).
-Summaries generated by LLMs showcase high coherence and are overwhelmingly preferred by the human evaluators (Pu et al., 2023; Liu et al., 2024).
-However, using LLMs for extractive summarization has two significant limitations: (i) as generative models, LLMs naturally tend to perform abstractive summarization by paraphrasing rather than selecting the most relevant sentences; and (ii) due to the finite size of the context window, LLMs cannot handle long texts in a single input, underscoring the need for a method that allows for processing long text.
-Given the increasing reliance on LLMs for summarization, it is worth investigating the utility of LLMs for extractive summarization of large volumes of user-generated text.
-To address these limitations, we present a novel framework LaMSUM (Large Language Model based Extractive SUMmarization) that integrates LLM-generated summaries with voting algorithms borrowed from Social Choice Theory (Brandt et al., 2016).
-Voting algorithms are traditionally employed in decision making processes to aggregate preferences and opinions of a large population; the prime example being the electoral system, where voters’ preferences for candidates are gathered to elect a winner that best represents the voters’ preferences.
-Voting systems can be categorized into single-winner and multi-winner types, depending on the number of winners.
-In a single-winner system, only one candidate is elected, whereas a multi-winner system allows for the selection of multiple candidates.
-Extractive summarization can be viewed as a multi-winner election, where the input units act as candidates, and the units chosen for the summary are considered as the winners.
-A multi-winner voting system includes approval-based committee (ABC) voting, where voters approve a subset of candidates without ranking, and ranked choice voting, where voters rank candidates by preference.
-In LaMSUM, we utilize two approval based voting algorithms – Plurality Voting (Mudambi et al., 1996) and Proportional Voting (Lackner et al., 2023), and one ranked based voting algorithm – Borda Count (Emerson, 2013).
+Extensive evaluation using three popular LLMs (Llama, Mistral and GPT-4o) demonstrates that LaMSUM outperforms state-of-the-art extractive summarization methods for Safe City posts.
+Overall, this work represents one of the first attempts to achieve extractive summarization through LLMs, and is likely to support stakeholders by offering a comprehensive overview and enabling them to develop effective policies to minimize incidents of unwarranted harassment.
+Warning: This paper contains content that may be disturbing or upsetting.
+Introduction
+  Category Post
+  Robbery This incident took place in the evening.Two bikers came on a bike and snatched Rs.17000 from an old lady at gun point.
+Stalking I was stalked by a guy who followed me for days and also he sent me letters on my doorstep saying he was madly in love with me and he is obsessed with my body.
+Sexual Invites I was walking on footpath to a place nearby to meet my friends.
+A man was driving his car on parallel road and was continuously passing vulgar comments.
+I ignored but 15 min later he stopped his car and said “chalri h ky, paise h mere pas” (I have money, you want to come with me?).
+I screamed and asked for help from people around me.
+Mas*****tion in public One day I reached my school prior to the school timings mistakenly. One of the drivers called me and my friend and started mas*****ting in front of us.
+Ogling I was going to my coaching by driving my scooty when suddenly some boys came on a bike.
+They started making cheap comments and teasing me.
+It was late and the area was secluded.
+I got scared as they started revolving their bike around me.
+I started driving in the direction of a crowded place that is when they left.
+Showing Po**ography A man in a car parked outside was watching po** when I saw him. He turned his device towards me and did offensive hand gestures inviting me.
+Sexual Assault When I was 7 years old, the shopkeeper removed my clothes and started touching me everywhere. He also tried to do it to another girl and failed.
+Domestic       Violence My husband always doubts on my character and doesn’t allow me to go outside alone, uses very vulgar language and beats me.
+Table 1: Examples of harassment cases shared on the Safe City platform.
+Proactive action by authorities and citizens can help prevent numerous such incidents.
+Providing stakeholders with a concise overview of incidents occurring in a specific area is crucial and this can be effectively achieved by utilizing summarization algorithms.
+In recent decades, the widespread availability of the internet has provided seamless access to online platforms to millions of people.
+Governments worldwide are increasingly utilizing these platforms to gather information directly from citizens – referred as Citizen Reporting (Kopackova and Libalova 2019).
+By leveraging tools such as mobile applications, web-based portals, and social media integrations, citizen reporting platforms establish a direct and efficient communication link between individuals and the relevant authorities, enabling faster issue resolution and facilitating active public participation in community improvement.
+Beyond immediate problem solving, real-time data gathered through these platforms contribute valuable information for urban planning and proactive measures, paving the way for more efficient and adaptive communities.
+Citizen reporting typically addresses topics such as community issues, environmental challenges, crime prevention, public health, and disaster response (Shin et al.
+2024).
+A notable citizen reporting platform is Safe City, developed by the nonprofit ‘Red Dot Foundation’ [^1] https://webapp.safecity.in; https://reddotfoundation.in, where people post incidents of sexual harassment, violence and assault.
+Table 1 showcases some example incidents shared by users on the platform.
+Although the platform is accessible worldwide via website and mobile app, the majority of its users are based in India.
+Despite making significant strides in recent decades in areas like Science & Technology, Defense and Agriculture, India continues to grapple with the critical issue of women safety.
+As per the ‘2023 Women Peace and Security Index’, India was positioned at 128th among 177 evaluated countries (Buchholz 2024).
+Violence against women remains an enduring and urgent concern, taking various forms, including domestic abuse, sexual harassment, rape, dowry-related violence, honor killings, and human trafficking.
+One such recent incident that deeply affected the nation was the rape and murder of a 31-year-old doctor while she was on duty in August 2024 (Times 2024).
+Several other publicized incidents include the 2020 Hathras case, where a young Dalit woman was brutally attacked and raped, ultimately resulting in her death (Dublish 2020) and the 2012 Nirbhaya incident involved the horrific gang rape and fatal assault of a 22-year-old woman in a private bus (Today 2020).
+While such horrific incidents cannot be entirely avoided through reporting alone, platforms like Safe City can play a crucial role in preventing certain cases of sexual assault.
+By enabling users to analyze reported incidents, assess the safety of specific locations, and make informed decisions when traveling to potential hotspots, these platforms contribute to enhanced personal safety and awareness.
+Similarly, the local authorities can also benefit from these platforms to assess emerging cases, identify the underlying factors and determine proactive measures for effective resolution.
+However, the challenge for the authorities is to navigate the high volume of information in such platforms.
+Manually reviewing all posts is often impractical, necessitating a need for a summarization algorithm that can identify and select posts that are diverse as well as representative of the original data.
+Additionally, platforms like Safe City often feature a curated selection of posts on their homepage to showcase their core purpose, mission, and services.
+This deliberate selection also acts as a form of summarization.
+Summarization algorithms are of two types: ‘extractive’ and ‘abstractive’.
+In extractive summarization, the algorithm selects a subset representative of the original text (Xu et al.
+2020; Zhong et al.
+2020; Zhang, Liu, and Zhang 2022, 2023a).
+In contrast, abstractive summarization algorithms generate summaries that capture the essence of the original text, often paraphrasing the content (Pu, Gao, and Wan 2023).
+For platforms like Safe City, extractive summarization is more suitable, as the goal is not to paraphrase the posts but to select a few that accurately capture a snapshot of the original content.
+When summarizing such sensitive posts, preserving the user’s exact words is essential, making extractive summarization particularly valuable in maintaining authenticity and context.
+Several extractive summarization algorithms for user generated content have been proposed in the literature, primarily for text written in English (Bhattacharya et al.
+2021; Kanwal and Rizzo 2022; Mukherjee et al.
+2020; Jia et al.
+2020).
+However, the Safe City platform receives posts in multiple languages from across India [^2] India is a linguistically rich and diverse country, home to 22 scheduled languages and many more spoken dialects., including numerous code-mixed entries where multiple languages are blended, such as Hinglish (a mix of Hindi and English).
+Such multilinguality limits the applicability of existing algorithms for extractive summarization of posts submitted to Safe City.
+In recent years, Large Language Models (LLMs) have demonstrated very good performance across various tasks in multilingual and code-mixed settings (Ouyang et al.
+2022; Brown et al.
+2020; Tang et al.
+2023a; Jin et al.
+2024b).
+Plus, summaries generated by LLMs showcase high coherence and are overwhelmingly preferred by human evaluators over other baseline algorithms (Pu, Gao, and Wan 2023; Liu et al.
+2024).
+These prior results motivated us to investigate the utility of LLMs for extractive summarization of large volumes of user generated posts.
+However, we encountered two significant limitations which hinder the immediate application of LLMs for extractive summarization: (i) as generative models, LLMs perform abstractive summarization by paraphrasing rather than selecting the most relevant sentences (as shown in Figure 1) (Worledge, Hashimoto, and Guestrin 2024); and (ii) due to the finite size of the context window, LLMs cannot handle long texts in a single input, underscoring the need for a method that allows for processing long text (Jin et al.
+2024a).
+To overcome these limitations, in this paper, we present a novel framework LaMSUM (Large Language Model based Extractive SUMmarization) that integrates LLM-generated summaries with voting algorithms borrowed from Social Choice Theory (Brandt et al.
+2016).
 Our judicial application of voting algorithms with a multi-level summarization framework ensures that LaMSUM outperforms the state-of-the-art fine-tuned summarization models.
 In summary, in this work, we make the following contributions:
-  • We propose a novel framework LaMSUM which can effectively summarize large (having ¿30K tokens) collection of user generated content.
-  • LaMSUM considers a multi-level summarization model that utilizes voting algorithms to combine outputs to generate robust summaries.
-  • Analyse whether an ensemble model with multiple LLMs in LaMSUM can lead to improved outcomes across different voting methods.
-To our knowledge, this is the first work to implement extractive summarization of large user-generated texts using LLMs by combining summarization with voting algorithms.
-We believe this work can spawn further research in this direction.
-2. Background and Related Work
+  • We propose a novel framework LaMSUM which can produce extractive summaries from large (having $>$30K tokens) collection of user generated content.
+LaMSUM considers a multi-level summarization model that utilizes voting algorithms to combine LLM outputs to generate robust summaries.
+  • Extensive experiments demonstrates that LaMSUM outperforms the state-of-the-art extractive summarization algorithms.
+  • We apply LaMSUM to user posts on the Safe City platform and develop a companion website that offers citizens and authorities a quick, comprehensive summary of harassment incidents occurring in localities across India.
+To our knowledge, this is the first work to implement extractive summarization of large collection of user-generated texts using LLMs by combining summarization with voting algorithms.
+At the same time, we demonstrate the effectiveness of such algorithms to facilitate data-driven decision-making promoting safer communities by providing actionable insights into reported incidents.
+We believe this work can spawn further research in this direction, and to enable it, we are making the dataset, including the human-annotated gold standard summaries, available upon request.
+Code is available at https://anonymous.4open.science/r/LaMSUM/
+Background and Related Work
 In this section, we review the relevant prior works that provide the foundation for our current research.
-Figure 1.
-Current LLMs, by default, produce abstractive summaries.
-Llama3-70b-8192, despite specifically prompted for extractive summarization, generates abstractive summaries.
+Figure 1: Current LLMs, by default, produce abstractive summaries.
+Llama-3.3-70b, despite specifically prompted for extractive summarization, generates abstractive summaries.
 This behavior underscores the need for a targeted approach to enable LLMs to effectively generate extractive summaries.
-Text Summarization Algorithms
-With the growing volume of online data, the demand for algorithms that automatically shorten and summarize texts is increasing.
-Automatic summarization can be approached in two ways: extractive and abstractive.
-In extractive summarization, a subset of input collection is selected for the summary based on their perceived quality and significance, aiming to represent a larger dataset with a concise sample.
-Abstractive summarization, on the other hand, generates natural language summaries that capture the most critical information from the original text.
-Both approaches seek to provide readers with a concise overview of the textual content.
-Over the years, many text summarization algorithms have been proposed in the literature; the reader can refer to (Gupta and Lehal, 2010; El-Kassas et al., 2021) for detailed surveys.
-Extractive Summarization of User Generated Content
-A large body of prior research has focused on summarizing long documents (Allahyari et al., 2017; Dong, 2018; Cajueiro et al., 2023).
-In fact, traditionally, summarization tasks have focused on summarizing a single document, such as a news article or a business report.
-In recent years, summarization has been increasingly applied on different types of user generated text (e.g., tweets, Facebook or Reddit posts) (Mackie et al., 2014; Inouye and Kalita, 2011; Jiang et al., 2021), where the task is to summarize short, independent posts written by many users.
-Several extractive summarization algorithms tailored specifically for user generated content have also been proposed (Sharifi et al., 2010; Xu et al., 2013; Dash et al., 2019b; Mukherjee et al., 2022; Rudra et al., 2018; Jones et al., 2023).
+AI Solution through Citizen Reporting
+Web and social media platforms receive posts on sensitive issues such as online harassment, hate speech, abusive behavior, violence etc.
+Abuse experienced by users lead to mental stress often forcing them to leave the platform (Sambasivan et al.
+2019; Thomas et al.
+2022; Kim et al.
+2024).
+Several AI-powered solutions have been designed to address solutions to these critical issues.
+Machine learning based classifiers and language models are utilized to detect the cases of sexual abuse, hate speech, offensive language, human trafficking and harassment cases (Sawhney et al.
+2021; Hassan et al.
+2020; Davidson et al.
+2017; Upadhayay, Lodhia, and Behzadan 2021; Stoop et al.
+2019; Ghosh Chowdhury et al.
+2019).
+Modelling the cyberbullying behavior using social network and language-based features can improve the classifier performance (Ziems, Vigfusson, and Morstatter 2020; Olteanu et al.
+2018).
+Development of a mobile computing based reporting tool empowers individuals with intellectual and developmental disabilities (I/DD) to self report abuse and share the incident with the intended group (Venkatasubramanian et al.
+2021; Sultana et al.
+2021).
+With the rise of public figures encouraging women to speak up, the number of non-anonymous self-reported assault stories have increased (ElSherief, Belding, and Nguyen 2017).
+Counterspeech is proven to be a viable alternative to blocking or suspending problematic messages or accounts, as it better aligns with the principles of free speech (Mathew et al.
+2019).
+Conversational agents (CAs) have attracted significant interest as potential counselors due to their features, such as anonymity, which can help address many challenges associated with human-human interaction (Park and Lee 2021).
 Large Language Models (LLMs) for Summarization
-LLMs are now being extensively used for summarization (Brown et al., 2020; Tang et al., 2023b; Jin et al., 2024).
-Multiple works have proposed few-shot learning frameworks for the abstractive summarization of news, documents, webpages, and generic texts (Zhang et al., 2023b; Tang et al., 2023a; Yang et al., 2023; Bražinskas et al., 2020; Laskar et al., 2023), but their primary focus remains on short documents that can fit in the LLM context window.
-Researchers have also observed that human evaluators are increasingly preferring LLM-generated summaries compared to other baselines (Zhang et al., 2024a; Wu et al., 2024; Goyal et al., 2023; Zhang et al., 2023c; Liu et al., 2024).
-Despite the advancements, recent studies have also uncovered factual inaccuracies and inconsistencies in LLM-generated summaries (Tang et al., 2024; Tam et al., 2023; Luo et al., 2023; Laban et al., 2023).
-Figure 2.
-LaMSUM: Multi-level framework for extractive summarization of large user-generated text.
+LLMs are now being extensively used for summarization (Brown et al.
+2020; Tang et al.
+2023a; Jin et al.
+2024b).
+Multiple works have proposed few-shot learning frameworks for the abstractive summarization of news, documents, webpages, and generic texts (Zhang, Liu, and Zhang 2023b; Tang et al.
+2023b; Yang et al.
+2023; Bražinskas, Lapata, and Titov 2020; Laskar et al.
+2023), but their primary focus remains on short documents that can fit in the LLM context window.
+Researchers have also observed that human evaluators are increasingly preferring LLM-generated summaries compared to other baselines (Zhang et al.
+2024; Wu et al.
+2024; Goyal, Li, and Durrett 2023; Zhang, Liu, and Zhang 2023c; Liu et al.
+2024).
+Despite the advancements, recent studies have also uncovered factual inaccuracies and inconsistencies in LLM-generated summaries (Tang et al.
+2024; Tam et al.
+2023; Luo, Xie, and Ananiadou 2023; Laban et al.
+2023).
+Figure 2: LaMSUM: Multi-level framework for extractive summarization of large user-generated text.
 Input set $\mathcal{T}$ (level 0) is divided into $\lceil{\frac{|\mathcal{T}|}{s}}\rceil$ chunks each of size s.
@@
 Iteratively the same procedure is repeated till we obtain a summary of size $k$.
-We set $q=k$ to ensure our algorithm can effectively handle the worst-case scenario where all the textual units in the final summary may come from the same input chunk (Section 3.2).
-Extractive Summarization through LLMs: The Current State By default, LLMs produce abstractive summaries, meaning that the summary text is distinct from the input text, even when it is instructed to do otherwise.
+We set $q=k$ to ensure our algorithm can effectively handle the worst-case scenario where all the textual units in the final summary may come from the same input chunk.
+Extractive Summarization through LLMs: The Current State
+By default, LLMs produce abstractive summaries, meaning that the summary text is distinct from the input text, even when it is instructed to do otherwise.
 To illustrate this, we present a small example in Figure 1.
-An LLM, when prompted, could clearly explain extractive summarization, yet, when we instructed it to perform extractive summarization on a set of 100 sentences, it fails to do so and instead generated an abstractive summary.
+An LLM, when prompted, could clearly explain extractive summarization, yet, when we instructed it to perform extractive summarization on a set of 50 sentences, it fails to do so and instead generated an abstractive summary.
 Prior to our current work, only two studies attempted to perform similar tasks.
-Zhang et al.
-(2023b) attempted summarization of short news articles using GPT 3.5, while Chang et al.
+Zhang, Liu, and Zhang (2023b) attempted summarization of short news articles using GPT 3.5, while Chang et al.
 (2024) attempted abstractive summarization for book-length documents.
-However, both these approaches suffer from practical limitations such as lack of contextual dependencies in user generated text and the problem with positional bias in LLMs.
+However, both these approaches suffer from practical limitations such as lack of contextual dependencies in user generated text and the problem with positional bias.
 To the best of our knowledge, ours is the first attempt to perform extractive summarization on a large collection of user generated texts through LLMs, while tackling the challenge of positional bias.
 We describe our proposal in detail in the next section.
-3. Generating Extractive Summaries through LLMs
+LaMSUM: Generating Extractive Summaries through LLMs
 In this section, we define the problem statement formally and introduce our novel summarization framework LaMSUM (Large Language Model based Extractive SUMmarization) that leverages LLMs to summarize large user-generated text.
-3.1. Task Formulation
+Task Formulation
 […]
-3.2. Multi-Level Summarization
+Multi-Level Summarization
 LLMs have a limited context window, making it impossible to input large text collections all at once.
 While recent models like GPT-4 support context windows of up to 128k tokens, they still cannot accommodate book-length inputs within a single window.
-Consequently, the input must be divided into smaller chunks to perform the desired task (Chang et al., 2024).
-Thus, LaMSUM employs a multi-level framework for extractive summarization, enabling it to consider input data of any size (detailed in Figure 2)  [^1] Code is available at https://anonymous.4open.science/r/LaMSUM/ .
+Consequently, the input must be divided into smaller chunks to perform the desired task (Chang et al.
+2024).
+Thus, LaMSUM employs a multi-level framework for extractive summarization, enabling it to consider input data of any size (detailed in Figure 2).
+.
 The set $\mathcal{T}$, which contains the original textual units, is provided as input at level 0 and is divided into $\lceil{\frac{|\mathcal{T}|}{s}}\rceil$ number of chunks of size $s$.
-From each chunk of size $s$, we generate a summary (discussed in Section 3.3) of size $q$ (where $q<s$), and repeat this process for all $\lceil{\frac{|\mathcal{T}|}{s}}\rceil$ chunks.
+From each chunk of size $s$, we generate a summary of size $q$ (where $q<s$), and repeat this process for all $\lceil{\frac{|\mathcal{T}|}{s}}\rceil$ chunks.
 [^3] Note that a chunk of size $s$ refers to a chunk containing $s$ textual units.
@@
 However, this approach assumes a uniform distribution of potential candidates across chunks that can be included in the final summary.
-In LaMSUM, we keep $q=k$ i.e., we extract $k$ textual units from each chunk eliminating the chance of missing any potential candidate (discussed in Section 4.4).
-In the worst-case scenario, all $k$ units in the final summary can come from a single chunk, and our algorithm can handle such cases effectively.
-It is important to note that we are dealing with user-generated content, such as tweets, which lack contextual connections.
-Unlike book summarization, where chapters are interconnected and the context of previous chapters is crucial for summarizing the current one, tweets are generally standalone and contextually independent.
+In LaMSUM, we keep $q=k$ i.e., we extract $k$ textual units from each chunk eliminating the chance of missing any potential candidate.
+In the worst-case scenario, all $k$ units in the final summary can come from a single chunk, and our algorithm can handle such cases effectively, as we keep $q=k$.
+It is important to note that we are dealing with user-generated content, such as posts, which lack contextual connections.
+Unlike book summarization, where chapters are interconnected and the context of previous chapters is crucial for summarizing the current one, posts are generally standalone and contextually independent.
 Thus, our approach of independently deriving summaries from each chunk works well in our setup, as each textual unit operates independently of the others and there are no long-range dependencies.
-Figure 3.
-Textual units (e.g., posts) in the input chunk are shuffled to account for the positional bias.
+Figure 3: Textual units (e.g., posts) in the input chunk are shuffled to account for the positional bias.
 $m$ different chunk variations are obtained through shuffling, which are subsequently summarized using LLMs.
 $m$ summaries are then aggregated by voting algorithms to get the final summary.
-3.3. Summarizing a Chunk
-Next, we discuss how LaMSUM summarizes a chunk (Algorithm 2) by tackling the positional bias in LLMs and leveraging voting algorithms drawn from Social Choice Theory (Brandt et al., 2016).
-3.3.1. Tackling Positional Bias
-Prior research  (Brown and Shokri, 2023; Zhang et al., 2023b; Jung et al., 2019; Wu et al., 2024) has highlighted that summarization using LLM is prone to positional bias, i.e., the sentences located in certain positions, such as the beginning of articles, are more likely to be considered in the summary.
+Summarizing a Chunk
+Next, we discuss how LaMSUM summarizes a chunk (Algorithm 2) by tackling the positional bias in LLMs and leveraging voting algorithms drawn from Social Choice Theory (Brandt et al. 2016).
+Tackling Positional Bias
+Prior research  (Brown and Shokri 2023; Zhang, Liu, and Zhang 2023b; Jung et al.
+2019; Wu et al.
+2024) has highlighted that summarization using LLM is prone to positional bias, i.e., the sentences located in certain positions, such as the beginning of articles, are more likely to be considered in the summary.
 To address this issue and generate a robust summary, we create $m$ different variations by shuffling the textual units within the input chunk.
 This ensures that each unit has the opportunity to appear in different positions within the input text (refer Figure 3).
-3.3.2. Zero-shot prompting
-For each input chunk, we obtain $m$ different summaries (one for each variation) by prompting the LLM. We employ the following two prompts to obtain the summaries –
-a
-Select the most suitable units that summarize the input text.
-Prompt: ”Input consists of ¡chunk_size¿ sentences.
-Each sentence is present in a new line.
-Each sentence contains a sentence number followed by text.
-You are an assistant that selects best ¡summary_length¿ sentences (subset) which summarizes the input.
-Think step by step and follow the instructions.
-¡sentences¿”
-b
-Generate a ranked list in descending order of preference.
-Prompt: ”Input consists of ¡chunk_size¿ sentences.
-Each sentence is present in a new line.
-Each sentence contains a sentence number followed by text.
-You are an assistant that outputs the sentences in the decreasing order of their relevance to be included in the summary.
-Think step by step and follow the instructions.
-¡sentences¿ Remember that output should contain all the sentences in the decreasing order of their preference.”
+Zero-shot prompting
+For each input chunk, we obtain $m$ different summaries (one for each variation) by prompting the LLM.
+We employ the following two prompt strategies to obtain the summaries (detailed in Appendix Table 7) –
+1
+Select most suitable units that summarize the input text.
+2 Generate a ranked list in descending order of preference.
 Algorithm 1 Algorithm for multi-level summarization
@@
          if $i=n_{chunks}$ then $\triangleright$ if last chunk
-              $ei=|\mathcal{T}|$triangleright ending index is equal to length of $\mathcal{T}$
+              $ei=|\mathcal{T}|$triangleright ei is equal to length of $\mathcal{T}$
          end if
-         $width=ei-si$triangleright number of textual units in a chunk
+         $width=ei-si$triangleright no. of textual units in a chunk
          if $width<=q$ then $\triangleright$ if last chunk
@@
      for $i\leftarrow 1$ to $m$ do $\triangleright$ for each variation of a chunk
-         $V=\textsc{Shuffle}(\mathcal{T},si,ei,i)$triangleright shuffle with random state $i$
-         $R=\textsc{LLM}(V,q)$triangleright obtain summary from the LLM
+         $V=\textsc{Shuffle}(\mathcal{T},si,ei,i)$triangleright shuffle with state $i$
+         $R=\textsc{LLM}(V,q)$triangleright obtain LLM summary
          $C=\textsc{Check}(R,\mathcal{T},si,ei)$triangleright output calibration
@@
      end for
-     return $\textsc{VOTING}(X,q)$triangleright voting to obtain the final summary
+     return $\textsc{VOTING}(X,q)$triangleright voting for final summary
 end function
   Original Post LLM Modified Output
-  My allergies be acting up &amp; Claritin be letting them &gt;,&lt; My allergies be acting up & Claritin be letting them.
-I spent the extra money to have fast relief, but instead I am SUFFERING #gimmemymoneyback I spent the extra money to have fast relief, but instead I am SUFFER
-Table 1. Examples illustrating that LLMs when selecting textual units for summarization, often demonstrate a propensity to alter certain words or introduce new ones.
-3.3.3. Output Calibration
+  In a train some people were staring me continuously. It was very uncomfortable Some people were staring me continuously in a train.
+We were going to metro station, a biker started following us. When we shouted, he rode away A biker followed us and rode away when we shouted.
+Some boy do dirty comments on me and my religion. Boys made dirty comments about my religion.
+Table 2: Examples illustrating that LLMs when selecting textual units for summarization, often demonstrate a propensity to alter certain words or introduce new ones.
+Output Calibration
 LLMs may alter certain words from the input text while generating extractive summaries, as shown in Table 2.
 Thus, we perform additional checks to ensure that the textual units selected in the summary are indeed a subset of $\mathcal{T}$.
-If the post selected by the LLM (say $x$) is not present in the original text $\mathcal{T}$, we identify the post with the closest resemblance to $x$ by computing the edit distance (Ristad and Yianilos, 1998).
+If the post selected by the LLM (say $x$) is not present in the original text $\mathcal{T}$, we identify the post with the closest resemblance to $x$ by computing the edit distance (Ristad and Yianilos 1998).
 LLMs may also hallucinate, generating new sentences rather than selecting units from the input.
-In such instances, the edit distance between the generated unit $x$ and all the original textual units would be high.
-We discard these sentences from the output.
-3.4. Reimagining Summarization as an Election
+In such cases, the edit distance between the generated unit $x$ and the original textual units tend to be high.
+To address this, we adopt an alternative approach – we extract key elements such as nouns, verbs, and adjectives from the newly generated sentence $x$.
+Then search for the exact same set of keywords with the textual units in $\mathcal{T}$ and identify the post with the highest number of matching keywords as the closest match to $x$ (refer Algorithm 4 in Appendix).
+Reimagining Summarization as an Election
 As mentioned earlier, for a given chunk, we obtain $m$ summaries – one for each variation.
 We imagine the process of creating the final summary from these $m$ summaries to be a multi-winner election, where the textual units in $m$ summaries correspond to ballots (candidates) and the role of the voting algorithm is to pick $q$ winners.
-We employ three different voting methods, namely Plurality Voting (Mudambi et al., 1996), Proportional Approval Voting (PAV) (Lackner et al., 2023) and Ranked Choice Voting (Emerson, 2013) to determine the final summary.
+We employ three different voting methods, namely Plurality Voting (Mudambi, Navarra, and Nicosia 1996), Proportional Approval Voting (PAV) (Lackner, Regner, and Krenn 2023) and Ranked Choice Voting (Emerson 2013) to determine the final summary.
 Due to the varying input requirements of different voting methods, changing both the prompting approach and the output generated by the LLM becomes imperative.
@@
 Since both plurality and proportional are approval-based voting algorithms, the units are either approved or disapproved by the underlying LLM, with no explicit ranking or preference order.
-In this case, we prompt the LLM to select the best ¡$q$¿ sentences that summarize the input text as shown in Section 3.3.2 a.
+In this case, we prompt the LLM to select the best $<q>$ sentences that summarize the input text as shown in 1.
 On the other hand, ranked choice voting entails assigning a score to each textual unit and subsequently selecting the highest-scoring units for inclusion in the summary.
-For ranked voting, we use the Borda count, a positional voting algorithm (Emerson, 2013).
+For ranked voting, we use the Borda count, a positional voting algorithm (Emerson 2013).
 In the Borda method, each candidate is assigned points corresponding to the number of candidates ranked below them: the lowest-ranked candidate receives 0 points, the next lowest gets 1 point, and so forth.
 The candidates with the highest aggregate points are declared as the winners.
-In ranked voting, we prompt the LLM to output sentences in descending order of their suitability for the summary as discussed in Section 3.3.2 b.
+In ranked voting, we prompt the LLM to output sentences in descending order of their suitability for the summary as discussed in 2.
 It is important to note that the prompting technique and the output generated by LLM vary for different voting methods.
@@
 Whereas in ranked choice voting, the output from LLM is a list of the same length as input i.e.
-$s$ with all the units sorted in decreasing order of their preference towards the summary, and Borda Count (Emerson, 2013) is used to identify the top $q$ textual units.
+$s$ with all the units sorted in decreasing order of their preference towards the summary, and Borda Count (Emerson 2013) is used to identify the top $q$ textual units.
 In the next section, we highlight how the voting-based summarization schemes outperform the Vanilla setup, which does not use voting.
-4. Experimental Setup
-      Parameters Claritin US-Election MeToo
-      #TextualUnits ($|\mathcal{T}|$) 3998 2107 483
-#InputWords 53609 35522 16737
-AV #CTU ($s$) 200 150 75
-#CSTU ($q$) 100 100 50
-RV #CTU ($s$) 40 40 40
-#CSTU ($q$) 20 20 20
-#SummaryTU ($k$) 100 100 50
-Table 2.
-Input parameters used for the proposed framework LaMSUM.
-#TextualUnits is the number of textual units in the input set i.e.
-$|\mathcal{T}|$.
-#InputWords represents the number of words present in the input set $\mathcal{T}$.
-#CTU is the number of textual units in a chunk i.e.
-chunk size represented as $s$.
-#CSTU is the number of textual units in the chunk summary represented as $q$.
-#SummaryTU is the number of textual units present in the final summary i.e.
-$k$.
-AV and RV represents Approval Voting and Ranked Voting respectively.
-4.1. Dataset
-Our experiments are conducted on three publicly available datasets, consisting of crowd-sourced data from X, listed in Table 2 (Dash et al., 2019a).
-The Claritin dataset contains 3,998 tweets about the benefits and the side-effects of the anti-allergic drug Claritin.
-US-Election dataset contains 2,107 tweets from 2016 US Presidential Election where people support and attack different political parties.
-Me-Too dataset includes 483 tweets from the October 2018 MeToo movement, where individuals recount the harassment cases they experienced.
-[^3] Dataset: https://github.com/ad93/FairSumm/tree/master/Dataset During pre-processing, we remove all web links and duplicate entries from the datasets.
-The datasets also include gold-standard summaries (reference summaries) which are human-generated i.e., the textual units that are strong candidates for inclusion in the summary are selected by the humans.
-Claritin and US-Election dataset, each have three gold standard summaries, comprising 100 textual units each.
-Me-Too dataset has two gold-standard summaries with 50 textual units each.
-4.2. Large Language Models (LLMs)
-LLMs are characterized by their extensive parameter sizes and remarkable learning abilities (Zhao et al., 2023; Chang et al., 2023).
-In our work, we utilize three open-source LLMs and one proprietary LLM to conduct experiments: llama3-70b-8192 from Meta (Touvron et al., 2023), mixtral-8x7b-32768 from Mistral AI (Jiang et al., 2024), gemini-1.0-pro from Google (Gemini Team, 2023) and
-gpt-4o-mini-2024-07-18 from OpenAI (OpenAI, 2024). Across all experiments, we keep temperature, top probability and output tokens as 0, 0.8 and 8192 respectively.
-4.3. Evaluation Metric
-For evaluating the quality of summaries generated by LaMSUM, we report ROUGE-1, ROUGE-2, and ROUGE-Lsum scores (Lin, 2004).
+Experimental Setup
+Dataset
+  No. Post Category A B C D E
+ PC1 Rape/Sexual Assault 33 27 17 23 18
+PC2 Chain Snatching/Robbery 49 16 103 32 22
+PC3 Domestic Violence 27 87 30 10 34
+PC4 Physical Assault 60 33 57 41 33
+PC5 Stalking 146 166 165 100 153
+PC6 Ogling/Staring 147 100 202 133 209
+PC7 Taking Photos 73 70 43 37 55
+PC8 Mas*****tion in public 45 55 27 30 42
+PC9 Touching/Groping 152 209 206 159 230
+PC10 Showing Po**ography 23 10 15 19 15
+PC11 Commenting/Sexual Invites 133 192 273 133 131
+PC12 Online Harassment 66 43 25 61 33
+PC13 Human Trafficking 3 4 2 2 1
+PC14 Others 50 23 36 41 16
+Table 3: The distribution of posts across each category for five datasets – City A, City B, City C, City D, and City E.
+Posts on the Safe City platform are tagged with various categories by their authors, with each category representing a different form of sexual harassment.
+                                                  Parameters City A City B City C City D City E
+                                                  Total no. of textual units ($|\mathcal{T}|$) 625 867 866 545 728
+Total no. of words in input set $\mathcal{T}$ 20544 12665 23501 22471 20807
+Approval Voting No. of textual units in a chunk ($s$) 100 100 100 100 100
+No. of textual units in chunk summary ($q$) 50 50 50 50 50
+Ranked Voting No. of textual units in a chunk ($s$) 40 40 40 40 40
+No. of textual units in chunk summary ($q$) 20 20 20 20 20
+No. of textual units in final summary ($k$) 50 50 50 50 50
+Table 4: Input parameters used for the proposed framework LaMSUM.
+$|\mathcal{T}|$ is the number of textual units in the input set.
+$s$ is the chunk size i.e.
+the number of textual units in a chunk.
+$q$ is the number of textual units in the chunk summary.
+$k$ is the number of textual units present in the final summary.
+We collect the Safe City posts from five different Indian cities [^4] The Red Dot Foundation has kindly made the data available for research purposes., which we refer to as City A, City B, City C, City D and City E.
+For City C and City E we obtain the posts for 3 years, i.e., from Dec 2021 to Nov 2024.
+For City A, City B and City D we consider the posts for 5 years, from Dec 2019 to Nov 2024.
+The rationale behind varying the duration of post selection is to keep the total number of posts below 1000, facilitating more efficient and accurate annotation by the human summarizers.
+Our dataset of City A, City B, City C, City D and City E consist of 625, 867, 866, 545 and 728 posts respectively.
+Sexual harassment can have various categories, such as physical assault, touching, stalking etc.
+Each post in the dataset is tagged with one or more categories by the author of the post.
+The distribution of posts across these categories is shown in Table 3.
+While the Safe City platform does not collect personal information such as names or identities, it does gather age, gender, and details of the incidents.
+Table 8 (in the Appendix) provides an overview of the attributes associated with each post.
+For our task, we focus solely on the main description provided in the posts.
+For each of the five city-specific datasets, we generate gold-standard (reference) summaries created by three domain experts.
+These experts carefully selected textual units from the posts that are strong candidates for inclusion in the summary.
+As a result, each dataset has three expert-generated gold-standard summaries, each comprising 50 textual units (posts).
+The expert annotators were provided with clear guidelines for selecting the posts for the reference summary – i) Diversity: Prioritize diverse posts that represent various forms of assault.
+ii) Descriptive: Give preference to posts with detailed descriptions over those containing only 2-3 words.
+iii) Severity: Include posts that depict more serious cases or require urgent attention.
+iv) Redundancy: Exclude posts that are repetitive or redundant.
+Additional details about the posts selected by the annotators is available in Figure 7 and Table 10 in the Appendix .
+Large Language Models (LLMs)
+LLMs are characterized by their extensive parameter sizes and remarkable learning abilities (Zhao et al.
+2023; Chang et al.
+2023).
+In our work, we utilize two open-source LLMs and one proprietary LLM to conduct experiments: llama-3.1-8B-instruct from Meta (Touvron and et al.
+2023), open-mistral-nemo-2407 from Mistral AI (Jiang and et al.
+2024) and gpt-4o-mini-2024-07-18 from OpenAI (OpenAI 2024).
+Across all experiments, we keep temperature, top probability and output tokens as 0, 0.8 and 8192 respectively.
+We downloaded llama-3.1-8B-instruct from Hugging Face and executed it locally on A6000 GPU.
+For open-mistral-nemo and gpt-4o-mini, we utilised official APIs.
+Evaluation Metric
+For evaluating the quality of summaries generated by LaMSUM, we report ROUGE-1, ROUGE-2, ROUGE-Lsum (Lin 2004).
 ROUGE-1, ROUGE-2 and ROUGE-L respectively evaluate the overlap of unigrams, bigrams and longest common subsequence between the generated summary and the reference summary.
 ROUGE-Lsum is more suitable for extractive summarization, as it applies ROUGE-L at sentence level and then aggregates all the results to obtain the final score.
-4.4. LaMSUM Input Parameters
+LaMSUM Input Parameters
 Input parameters for LaMSUM (Algorithm 1), such as $|\mathcal{T}|$ (total number of textual units in the set), $s$ (chunk size) and $k$ (length of summary) are listed in Table 4 for different voting algorithms and datasets.
@@
 The optimal value of $q$ that can handle worst case and also reduce the number of levels in multi-level summarization is $k$, thus we keep $q=k$ for experiments with approval voting.
-For instance, if $s$ is 200 and $q$ is 100, this indicates that only 50% of the units from each chunk advance to the next level.
+For instance, if $s$ is 100 and $q$ is 50, this indicates that only 50% of the units from each chunk advance to the next level.
 In ranked voting algorithm, we maintain smaller value for chunk size ($s$) to ensure that the LLMs output, which is of size $s$, fits within the context window.
@@
 For ranked voting, we set $s$ and $q$ to 40 and 20 respectively, upholding the selection ratio of 50% at each level.
-5. Experimental Evaluation
+Experimental Evaluation
 In this section, we present the empirical comparison of LaMSUM with competent baseline models and voting algorithms across datasets (refer Table 5).
-                      Models Claritin US-Election         MeToo
-R1 R2 RLSum R1 R2 RLSum R1 R2 RLSum
-                     ClusterRank 50.08 14.17 48.64 55.18 13.09 53.51 55.81 55.81 52.48
-DSDR 28.01 8.40 27.45 48.04 9.53 46.44 55.03 17.30 52.04
-LexRank 45.04 19.71 44.74 42.63 10.78 41.64 42.70 11.32 40.91
-SummBasic 58.25 19.29 56.76 55.36 12.43 53.94 58.23 21.53 56.07
-GPT2 61.61 23.58 60.74 55.86 15.07 54.81 40.63 11.24 38.86
-BERT 57.30 22.37 56.21 55.89 15.44 55.00 45.72 10.76 43.50
-XLNET 55.52 21.37 54.75 56.48 15.72 55.41 36.58 08.50 34.48
-BERTSUM (Liu and Lapata, 2019) 57.87 22.75 55.96 59.00 17.51 57.41 57.11 23.08 54.84
- Vanilla Llama3 58.07 21.86 56.73 56.12 12.93 54.71 51.84 17.84 49.47
-Mixtral 57.80 20.14 55.81 57.11 14.20 55.79 49.63 11.29 46.16
-Gemini 49.70 18.33 48.78 51.00 12.36 49.19 45.51 8.95 42.51
-GPT-4o mini 62.93 24.82 61.08 55.06 15.01 53.88 36.30 7.05 34.54
-  Plurality Voting Llama3 61.28 23.79 59.54 60.11 18.26 58.99 55.51 18.79 52.85
-Mixtral 59.13 22.90 57.40 59.55 16.01 58.12 55.41 17.62 52.63
-Gemini 55.43 19.32 53.35 58.83 16.25 57.54 57.95 21.47 54.87
-GPT-4o mini 64.20 26.71 62.66 58.15 15.91 56.78 54.55 19.66 51.9
-Proportional Voting Llama3 61.31 23.36 59.73 58.90 15.99 57.55 58.14 14.99 55.46
-Mixtral 60.30 24.03 58.58 58.46 15.74 57.28 57.36 22.97 54.82
-Gemini 61.77 19.83 59.81 57.92 15.85 56.49 54.02 13.95 50.90
-GPT-4o mini 64.13 26.75 62.30 58.68 19.08 57.62 48.55 13.23 45.35
- Borda Count Llama3 51.87 15.93 49.61 51.95 13.49 50.60 53.35 16.99 50.38
-Mixtral 56.07 22.28 54.65 48.71 13.37 47.48 53.24 18.80 50.96
-Gemini 51.45 16.92 49.39 54.01 13.38 52.56 50.13 21.45 47.96
-GPT-4o mini 58.01 18.08 56.05 53.10 15.94 51.63 50.10 17.82 47.75
-Table 3.
-Table showing metric scores from different models for various datasets.
+  City A City B City C City D City E
+Models R1 R2 RLSum R1 R2 RLSum R1 R2 RLSum R1 R2 RLSum R1 R2 RLSum
+ LexRank 57.644 28.084 55.904 38.898 16.909 38.239 56.379 24.897 54.057 46.221 21.709 44.612 52.157 23.681 49.812
+SummBasic 52.498 21.848 50.700 26.714 9.536 26.283 53.172 21.600 51.136 43.215 19.841 41.392 50.677 21.357 48.482
+LSA 55.250 25.156 53.219 52.529 25.947 51.462 45.623 17.378 43.981 42.370 18.262 40.762 47.522 17.521 45.585
+BERT 55.645 23.673 54.236 53.799 26.428 52.711 53.610 22.353 51.982 43.986 19.738 42.322 50.444 21.470 49.377
+XLNET 50.559 19.373 48.959 51.706 24.259 50.937 52.819 21.137 51.189 43.768 22.133 42.391 51.650 22.151 50.142
+BERTSUM 56.570 25.323 54.226 51.137 24.551 49.386 53.927 22.647 50.993 53.767 25.846 49.376 47.735 20.606 45.295
+LaMSUM 60.440 30.006 58.355 56.296 32.109 55.229 58.708 29.022 56.428 54.855 28.124 52.504 55.921 29.919 53.426
+Table 5: Table showing metric scores from different models for various datasets.
 Here, R1 = ROUGE-1 Score, R2 = ROUGE-2 Score, RLSum = ROUGE-LSum Score.
-The best value per dataset is shown in bold and clearly Approval voting outperforms all the other methods across all the evaluation measures.
-5.1. Baseline Comparison
-We compare LaMSUM with the pre-neural models (ClusterRank (Garg et al., 2009), DSDR (He et al., 2012), LexRank (Erkan and Radev, 2004), SummBasic (Nenkova and Vanderwende, 2005)), transformer based models (GPT2 (Radford et al., 2019), BERT (Miller, 2019), XLNET (Yang et al., 2019)) and with fine-tuned BERTSUM (Liu and Lapata, 2019) model.
-As presented in Table 3 it is observed that LaMSUM outperform the state-of-the-art summarization models.
-Earlier work (Zhang et al., 2023b) reported that the ChatGPT model achieves lower ROUGE scores on CNN/DM and XSum dataset.
-But our results demonstrate that our proposed framework LaMSUM performs significantly better than other fine-tuned models for large user-generated text.
-Results from the LLM exhibit variability when executed multiple times over the same input.
-We conducted experiments with LLMs for five iterations for the same input.
-Table 3 contains the maximum ROUGE score obtained from these five iterations and Table 6 (in supplementary material) displays the variance in the ROUGE scores when executed for five times.
+The best result for each dataset is shown in bold and clearly LaMSUM outperforms all the other methods across all the evaluation measures.
+Baseline Comparison
+We compare LaMSUM with the pre-neural models such as (LexRank (Erkan and Radev 2004), SummBasic (Nenkova and Vanderwende 2005), LSA (Gong and Liu 2001)), transformer based models (BERT (Miller 2019), XLNET (Yang et al.
+2019)) and with state-of-the-art fine-tuned BERTSUM (Liu and Lapata 2019) model.
+Our proposed method, LaMSUM, achieves optimal performance when gpt-4o-mini is used with Proportional Approval Voting.
+Therefore, we present the results of this combination as the outcomes of LaMSUM.
+As shown in Table 5, it is observed that LaMSUM surpasses state-of-the-art summarization models across all metrics.
+Does LaMSUM perform better than Vanilla LLM?
+Figure 4: Metric scores obtained through four different LLM setups.
+(i) Vanilla LLM without shuffling and voting method (ii) LaMSUM with Plurality Approval Voting (iii) LaMSUM with Proportional Approval Voting and (iv) LaMSUM with Borda Count Ranked Voting.
+Results demonstrate that gpt-4o-mini with Proportional Approval Voting performs the best across all the cases.
+Here, Llama, Mixtral and GPT refers to llama-3.1-8B, open-mistral-nemo, and gpt-4o-mini respectively.
+Our proposed framework, LaMSUM, ensures robust summary generation by shuffling and employing a voting algorithm to select the best textual units for the summary.
+It is crucial to compare LaMSUM with a multi-level LLM that does not use shuffling and voting, which we call Vanilla LLM.
+Algorithm 3 outlines the steps used by vanilla LLM to find the chunk summary.
+Figure 4 demonstrates that the vanilla multi-level LLM has lower ROUGE scores for each LLM compared to the proposed framework LaMSUM, indicating that shuffling and voting enhances the performance.
+Earlier work (Zhang, Liu, and Zhang 2023b) reported that the ChatGPT model achieves lower ROUGE scores on CNN/DM and XSum dataset.
+But our results demonstrate that our proposed framework performs significantly better than other fine-tuned language models such as BERTSUM for large user-generated text.
+Which LLM takes the lead?
+We conducted experiments using three LLMs – llama-3.1-8B, open-mistral-nemo, and gpt-4o-mini.
+As shown in Figure 4, gpt-4o-mini demonstrated the best performance among the three.
+While we utilized smaller models for these experiments, leveraging larger LLMs with higher parameter counts is likely to result in even better performance.
+Which voting algorithm performs the best?
+We experimented with three voting algorithms, two approval-based (pluarlity and proportional) and one ranked-based (borda-count).
+Experimental results indicate that LLMs with proportional approval voting perform the best compared to the other voting algorithms.
+As shown in Figure 4, it can be observed that PAV performs the best across all the LLMs.
+We hypothesized that rank-based voting would yield better results, as it makes more informed decisions about the potential sentences to be included in the summary.
+Contrary to our expectations, rank-based algorithms did not surpass the proportional approval voting.
+This can be attributed to multiple factors: (i) LLMs may hallucinate and output sentences in the same or in the reverse order as they were in the input.
+(ii) Occasionally, LLMs do not output all the sentences from the input, resulting in the padding of left-out sentences towards the end of the list, which disturbs the ranking and potentially affects the result.
+To overcome these problems, we kept the chunk size low, but the results still did not surpass those of the proportional approval-based voting algorithm.
 Algorithm 3 Algorithm for summarization of a chunk in Vanilla LLM
 function ChunkResult($\mathcal{T},si,ei,q,m$)
-     $R=\textsc{LLM}(\mathcal{T},si,ei,q)$triangleright $q$ textual units from $\mathcal{T}\in[si,ei]$
+     $R=\textsc{LLM}(\mathcal{T},si,ei,q)$triangleright $q$ textual units from $[si,ei]$
      $C=\textsc{Check}(R,\mathcal{T},si,ei)$triangleright output calibration
@@
 end function
-5.2. Vanilla LLM vs. LaMSUM
-Our proposed framework, LaMSUM, ensures robust summary generation by shuffling and employing a voting algorithm to select the best textual units for the summary.
-It is crucial to compare LaMSUM with a multi-level LLM that does not use shuffling and voting, which we call Vanilla LLM.
-Algorithm 3 outlines the steps used by vanilla LLM to find the chunk summary.
-Table 3 demonstrates that the vanilla multi-level LLM has lower ROUGE scores for each LLM compared to the proposed framework LaMSUM, indicating that shuffling and voting enhances the performance.
-5.3. Which Voting Algorithm Holds the Lead?
-We experimented with three voting algorithms, two approval-based and one ranked-based.
-Experimental results indicate that LLMs with approval voting perform the best compared to the ranked voting algorithm.
-We hypothesized that rank-based voting would yield better results, as it makes more informed decisions about the potential sentences to be included in the summary.
-Contrary to our expectations, rank-based algorithms performed even worse than neural and transformer-based models.
-This can be attributed to multiple factors: (i) LLMs hallucinate and output sentences in the same or in the reverse order as they were in the input.
-(ii) Occasionally, LLMs do not output all the sentences from the input, resulting in the padding of left-out sentences towards the end of the list, which disturbs the ranking and potentially affects the result.
-To overcome these problems, we kept the chunk size low as discussed in Section 4.4, but the results still did not surpass those of the approval-based voting algorithm.
-Takeaway: LLMs, when prompted to select sentences that can summarize the input, perform better than when tasked to rank the sentences in the order of their preference towards the summary.
-5.4. What Fails to Deliver Results?
+Analyzing the LaMSUM output
+Figure 5: Posts chosen by LaMSUM tend to be detailed and descriptive, offering a deeper level of information.
+Number of words in LaMSUM selected posts is often highest across various datasets, ensuring extensive and comprehensive summarization.
+  Models City A City B City C City D City E
+ LexRank 7.486 5.819 7.637 7.548 7.481
+SummBasic 8.198 5.734 8.050 8.020 8.196
+LSA 8.251 7.061 8.481 8.194 8.387
+BERT 8.068 7.762 8.437 8.186 8.191
+XLNET 8.144 7.689 8.619 8.205 8.400
+BERTSUM 8.331 7.539 8.690 7.960 8.112
+LaMSUM 8.563 7.822 8.622 8.606 8.480
+Table 6: Entropy values representing the diversity in the summaries produced by various algorithms.
+Bold values highlight the best and the underline denotes the second-best performance.
+LaMSUM achieves the highest diversity score across four datasets.
+We analyse the difference between the posts selected by the LaMSUM and those chosen by other algorithms.
+LaMSUM selects the post that are more descriptive and rich in detail, in contrast to posts that lack sufficient information.
+As shown in Figure 5, it is evident that across nearly all datasets, the posts selected by LaMSUM exhibit a higher word count.
+This indicates that the proposed algorithm is capable of capturing more detailed information compared to the others algorithms (refer Table 11 in Appendix for examples).
+Furthermore, posts selected by LaMSUM exhibit greater diversity, encompassing a broader range of harassment categories compared to the other baselines.
+We use entropy as a measure of diversity where a higher value is indicative of more randomness (Jost 2006).
+Table 6 demonstrates that LaMSUM generates diverse summaries as compared to other baselines.
+Accompanying Website
+We have demonstrated so far that LaMSUM generates superior summaries for Safe City posts.
+Building on this, we apply LaMSUM to posts from various localities and build a companion website for Safe City.
+This website enables the end-users and authorities to quickly access an overview of incidents in a specific area and customized time frame, enhancing awareness and helping policy action.
+For simplicity, we segregate the posts selected in the summary on the basis of the post categories.
+Figure 6 presents a screenshot of our accompanying website, which aims to assist users in making more informed decisions.
+[^5] The website link is hidden to maintain anonymity.
+Figure 6: Homepage of the accompanying website which allows users to obtain a quick snapshot of the incidents happening in an area. Results can be customized as per the district and the timeframes.
+Concluding Discussion
+Safe City platform receives numerous posts related to sexual harassment, having a summarization algorithm enables end-users to quickly review significant posts.
+This work marks an early attempt to achieve extractive summarization of large user-generated text that exceeds a single context window using zero-shot learning.
+The proposed multi-level framework LaMSUM leverages approval based and ranked based voting algorithms to generate robust summaries.
+Experiments conducted on Safe City crowd-sourced dataset demonstrated the efficacy of LaMSUM, as it outperformed the results achieved by state-of-the-art models.
+Note that there can be a concern regarding the potential data leakage, as the experiments involve newer LLMs that may have been exposed to the experimented datasets during their pre-training phase.
+We showcased that the vanilla LLM, which also includes LLM underperformed, whereas our proposed framework which generates robust summaries yielded good results.
+This highlights the efficacy of our model, even when it is exposed to data leakage.
+Limitation.
+Our proposed framework, LaMSUM, very well handles text of any length, conditioned on the fact that the final summary fits within a single context window.
+Some modifications to LaMSUM may be necessary when the output summary exceeds the size of a single context window.
+Ethical Considerations.
+Our research focuses on using LLMs to produce extractive summaries for platforms like Safe City.
+LLMs often exhibit bias towards their training data, which can influence their preference for certain textual elements during the summarization process.
+Their “black box” nature, with an opaque decision-making process, makes it challenging to discern how or why specific textual units are chosen for summarization.
+The posts selected by LaMSUM may not accurately represent real-life scenarios and cannot serve as a reliable proxy for actual situations.
+Additionally, LaMSUM may overlook less frequent posts with limited informational content.
+As a result, exclusive reliance on our framework could lead to the oversight of specific issues by the authorities.
+While LaMSUM can produce high-quality summaries, its use must be approached with careful consideration of potential ethical implications.
+Appendix A Appendix
+Algorithm
+Algorithm 4 shows the method for output calibration by utilizing two modules – i). minimum edit distance and ii). maximum count of keywords.
+Algorithm 4 Algorithm for output calibration
+function Check($R,\mathcal{T},si,ei$)
+     $Y=\{\}$triangleright store the result
+     for $x$ in $R$ do $\triangleright$ for each sentence $x$ in LLM result $R$
+         $min\_dist=\infty$triangleright keep track of min distance
+         $min\_idx=-1$triangleright post with min edit distance
+         $max\_count=0$triangleright matching keywords
+         $max\_idx=-1$triangleright post with max keywords
+         $K=\textsc{Keywords}(x)$triangleright obtain keywords in $x$
+         for $i\leftarrow si$ to $ei$ do $\triangleright$ for each unit in $\mathcal{T}$
+              $d=\textsc{EditDist}(x,\mathcal{T}[i])$triangleright obtain edit distance
+              if $d<min\_dist$ then $\triangleright$ lesser edit distance
+                  $min\_dist=d$triangleright update $min\_dist$
+                  $min\_idx=i$triangleright update $min\_idx$
+              end if
+              $c=\textsc{Count}(K,\mathcal{T}[i])$triangleright # keywords in $\mathcal{T}[i]$
+              if $c>max\_count$ then $\triangleright$ lesser edit distance
+                  $max\_count=c$triangleright update $max\_count$
+                  $max\_idx=i$triangleright update $max\_idx$
+              end if
+         end for
+         if $min\_dist<\epsilon$ then $\triangleright$ edit distance is low
+              $Y.add(\mathcal{T}[min\_idx])$triangleright add to result $Y$
+         else
+              $Y.add(\mathcal{T}[max\_idx])$triangleright add to result $Y$
+         end if
+     end for
+     return $Y$
+end function
+Zero Shot Prompting
+  Prompts
+ 1 Select the most suitable units that summarize the input text.
+Prompt: Input consists of $<$chunk_size$>$ sentences.
+Each sentence is present in a new line.
+Each sentence contains a sentence number followed by text.
+You are an assistant that selects best $<$summary_length$>$ sentences (subset) which summarizes the input.
+Think step by step and follow the instructions.
+$<$sentences$>$
+ 2 Generate a ranked list in descending order of preference.
+Prompt: Input consists of $<$chunk_size$>$ sentences.
+Each sentence is present in a new line.
+Each sentence contains a sentence number followed by text.
+You are an assistant that selects best $<$summary_length$>$ sentences (subset) which summarizes the input.
+Think step by step and follow the instructions.
+$<$sentences$>$
+Table 7: Prompts utilised for 1 Approval and 2 Ranked based voting algorithm.
+What Fails to Deliver Extractive Summary ?
 To ensure extractive summarization, we tested an additional approach – each sentence is tagged with a sentence number, LLM is prompted to select the best $q$ sentences and output only the sentence numbers of the best $q$ sentences.
 Thereafter, the sentences corresponding to the sentence numbers can be retrieved.
-For instance, if $s$ is 200 and $q$ is 100, the task is to output the sentence numbers of the best 100 sentences from a pool of 200 sentences.
+For instance, if $s$ is 100 and $q$ is 50, the task is to output the sentence numbers of the best 50 sentences from a pool of 100 sentences.
 In such cases, LLMs hallucinate and provide an output consisting of either all the odd number sentences or all the even number sentences.
 Takeaway: For extractive summarization, relying solely on indexes may result in hallucination, underscoring the importance of emitting the input content and not the numbers.
-6. Ensembling of LLMs in LaMSUM
-Figure 4.
-Ensembling applied at chunk using all the four models.
-$q$ length summaries are obtained from each chunk through different LLMs using the shuffling method discussed in Section 3.3.
-All the summaries are ensembled through voting algorithms to achieve the final summary of length $q$.
-This final $q$ length summary is thereafter forwarded to the next level.
- Ensemble Voting Method R1 R2 RLSum
- All 4 models Plurality 54.16 20.79 53.00
-Proportional 54.70 19.95 53.14
-Ranked 51.78 16.97 50.24
-Best 3 models Plurality 58.88 21.74 57.60
-Proportional 58.04 22.78 56.71
-Ranked 57.22 19.51 55.33
-Weak 3 models Plurality 50.64 18.82 49.31
-Proportional 52.92 19.23 51.07
-Ranked 47.50 16.74 46.37
-Table 4.
-Results obtained through ensembling of LLMs in LaMSUM for Claritin dataset.
-We compare 3 different cases: i) ensemble of all 4 models – Llama, Mixtral, Gemini and GPT-4o mini ii).
-ensemble of best 3 models – Llama, Mixtral and GPT-4o mini iii).
-ensemble of weak 3 models – Llama, Mixtral and Gemini.
-Different LLMs can be considered as distinct experts, each offering their vote or opinion about the textual unit to be included in the summary.
-In LaMSUM, each expert (LLM) operated independently, determining which sentences were worthy of inclusion.
-We hypothesize that using an ensemble of LLMs in LaMSUM, where these experts collaborate and reach a consensus, has potential to improve performance of the overall system (Verga et al., 2024; Schoenegger et al., 2024).
-We apply ensembling at each chunk as shown in Figure 4.
-For a given chunk $i$ containing $s$ textual units, the goal is to produce a final summary of size $q$ by gathering opinions of all the experts.
-Ensembling takes place in two stages.
-Stage 1 - For each chunk we obtain $q$ length summary through different LLMs by using shuffling and voting method as demonstrated in Section 3.3.
-Stage 2 - To ensemble the $q$ length summaries from different LLMs, we again leverage the voting methods.
-In stage 2 we make use of the same voting method that was used in stage 1 to maintain consistency.
-The summary generated from the ensembling process is treated as the final summary for chunk $i$ and is carried forward to the next level.
-As shown in Table 3, for Claritin dataset, GPT-4o mini outperforms the other models, while Gemini delivers weaker results.
-We evaluate three different scenarios based on the number of LLMs used in the ensembling process: i) all four models – Llama, Mixtral, Gemini and GPT-4o mini, ii) the top three models – Llama, Mixtral and GPT-4o mini, iii) the three weaker models – Llama, Mixtral and Gemini.
-Table 4 shows the results of ensembling in LaMSUM for different scenarios and voting methods.
-In the approval-based voting algorithm, using all four models or just the top three models for ensembling doesn’t significantly improve performance; the resulting performance after ensembling remains close to that of the weakest model in each case.
-However, ensembling with the ranked-based voting algorithm yielded interesting results.
-When ranked voting is applied to individual LLMs, it delivers the worst performance compared to other voting algorithms, as shown in Table 3.
-When using ranked voting with all four models in the ensembled LaMSUM, the performance is close to that of the weakest model – Gemini (49.39) ¡ Llama (49.61) ¡ Ensemble (50.24) ¡ Mixtral (54.65) ¡ GPT-4o mini (56.05).
-However, when considering the ensembled LaMSUM with the top three models – Gemini (49.39) ¡ Llama (49.61) ¡ Mixtral (54.65) ¡ Ensemble (55.33) ¡ GPT-4o mini(56.05), ensemble outperforms Mixtral, and the ROUGE score increases from 50.24 to 55.33.
-While the ensembled LaMSUM with the top three models does not surpass GPT-4o mini, it does enhance the performance of the overall setup.
-Takeaway: Ensemble of LLMs in LaMSUM yield good results with rank-based voting method, and can help improve the performance of the overall setup if we select models for ensembling that perform well individually in LaMSUM.
-Including all models in the ensemble could result in the weakest model becoming the bottleneck.
-7. Does Fine-Tuning Help?
-The performance of LLMs on specific downstream tasks, such as extractive summarization, can be suboptimal when not explicitly fine-tuned.
-LLMs tend to generate erroneous information while sounding persuasive and assured (Bang et al., 2023; Xu et al., 2024).
-In Section 2, we highlighted instances where LLMs exhibit nonplussed responses by generating abstractive summaries despite being prompted for extractive summarization.
-They also alter some words, so we perform output calibration in LaMSUM.
-To address these challenges, we explore the efficacy of instruction fine-tuning through explicit instructions (Wei et al., 2022) and compare it with LaMSUM.
-We utilize the Claritin dataset for instruction fine-tuning as it has the highest number of tweets among all the three datasets.
-It is divided into training (60%), validation (10%), and test (30%) set.
-The US-Election and MeToo datasets are retained as additional test sets for evaluation.
-Instruction fine-tuning involves three key components: i) Instruction: The prompts used for zero-shot prompting, as detailed in Section 3.3.2.
-ii) Input: Set of input text $\mathcal{T}$ to be summarized.
-iii) Output: The reference summary comprises sentences extracted from $\mathcal{T}$.
-We fine-tune an open-source LLM, llama3.1-8b from Meta (Dubey and et al., 2024), using a 15 GB Tesla T4 GPU.
-During the fine-tuning phase, we set the learning rate, warmup ratio, and batch size to 2e-4, 0.0, and 2, respectively.
-For this process, we employ PEFT (Parameter Efficient Fine Tuning), which involves freezing the layers of the pre-trained model and only fine-tuning the last few layers specific to the downstream task.
-We follow the QLoRA (Dettmers et al., 2023) method in PEFT training.
-QLoRA is an optimized variant of LoRA (Hu et al., 2021) that reduces the precision of weight parameters to 4-bit.
-This reduction in precision decreases the model size, which is advantageous in scenarios with limited memory available for fine-tuning.
-Table 5 shows the results of the fine-tuned llama3.1-8b model.
-We observed that the outputs are robust and adhere closely to the instructions provided.
-The generated summaries contain the exact sentences extracted from the input text, but the performance of the fine-tuned model is inferior compared to LaMSUM.
- Dataset R1 R2 RLSum
- Claritin 18.60 3.11 17.14
-US-Election 26.55 14.81 25.05
-MeToo 28.94 10.81 24.80
-Table 5. Metric scores from fine-tuned llama3.1-8b for various datasets.
-Takeaway: Fine-tuning an LLM with the appropriate instructions, input, and output can facilitate extractive summarization.
-But fine-tuning requires a substantial amount of data with gold-standard summaries.
-With less data, the model may overfit, resulting in poor performance, as shown in the Table 5.
-Additionally, a model fine-tuned on one dataset may not perform well on another dataset.
-Our proposed LaMSUM addresses these issues by being effective when the input data is large enough to exceed a single context window but small enough to be used for fine-tuning.
-8. Concluding Discussion
-This work marks an early attempt to achieve extractive summarization of large user-generated text that exceeds a single context window using zero-shot learning.
-The proposed multi-level framework LaMSUM leverages approval based and ranked based voting algorithms to generate robust summaries.
-Experiments conducted on three distinct crowd-sourced datasets demonstrated the efficacy of LaMSUM, as it outperformed the results achieved by state-of-the-art fine-tuned models.
-Ensembling of LLMs in our proposed framework can enhance the performance of weaker models.
-Note that there can be a concern regarding the potential data leakage, as the experiments involve newer LLMs that may have been exposed to the experimented datasets during their pre-training phase.
-In Section 5.2 we showcased that the vanilla LLM, which also includes LLM underperformed, whereas our proposed framework which generates robust summaries yielded good results.
-This highlights the efficacy of our model, even when it is exposed to data leakage.
-Limitation
-Our proposed framework, LaMSUM, very well handles text of any length, conditioned on the fact that the final summary fits within a single context window.
-Some modifications to LaMSUM may be necessary when the output summary exceeds the size of a single context window.
-We focused only on the user-generated text where each textual unit is independent of the others; future research could extend the framework to summarization tasks involving contextual dependency such as book summarization.
-Ethical Considerations
-Our research focuses on using LLMs to produce extractive summaries for user-generated text.
-Given the recent rise of LLMs and the growing interest in applying them across different research fields, we believe this research direction can help unveil the potential of LLMs.
-Below, we outline concerns that need to be considered and addressed in this research area:
-  • Bias and discrimination: LLMs are biased towards their training data, which can lead them to favor certain textual units when creating summaries.
-  • Transparency and accountability: LLMs are black boxes, with an opaque decision-making process, making it difficult to discern how or why specific textual units are chosen for summarization.
-This lack of transparency can create ethical issues, especially in scenarios where clear explanations for decisions are needed.
-  • Environmental impact: The training and operation of LLMs require substantial computational resources, which contribute to increased energy consumption and may have a negative environmental impact.
-While LLMs can produce high-quality summaries, their use must be approached with careful consideration of potential ethical implications.
+Safe City Post Features
+  Feature Details
+ id unique id for each post
+lang_id language id
+building building where incident took place
+landmark landmark near the place of incident
+area area where incident occurred
+city city where incident happened
+state name of the state
+country country name
+latitude coordinates information
+longitude coordinates information
+created_on date when the post is made
+description details about the incident
+additional_detail more information about the incident
+age age of the person
+gender_id gender id
+gender gender of the person
+incident_date date when the incident took place
+is_date_estimate binary value – yes or no
+time_from start time of the incident
+time_to end time of the incident
+is_time_estimate binary value – yes or no
+categories harassment category
+Table 8: Features or attributes associated with a post in Safe City platform. For our task, we utilise only the description feature.
+  City A City B City C City D City E
+Models R1 R2 RLSum R1 R2 RLSum R1 R2 RLSum R1 R2 RLSum R1 R2 RLSum
+                 Vanilla LLM
+ Llama 56.506 25.493 54.098 48.798 25.393 47.674 53.321 24.494 51.082 49.976 25.516 47.848 51.713 23.452 49.015
+Mixtral 53.009 21.994 49.848 42.948 19.178 41.970 52.337 24.547 49.392 48.066 26.513 46.000 52.250 24.946 50.195
+GPT 57.666 25.543 55.044 46.518 24.491 44.952 56.255 26.828 54.140 52.144 29.078 50.136 50.199 23.773 47.840
+                 LaMSUM + Plurality Voting
+ Llama 57.679 26.446 55.275 45.990 19.033 44.249 54.789 25.282 52.364 50.119 23.849 47.773 52.118 22.047 49.264
+Mixtral 56.757 24.740 53.805 43.639 18.116 41.414 53.511 23.155 50.916 52.277 26.102 48.121 53.538 25.938 50.562
+GPT 58.140 28.705 55.733 53.471 26.895 52.017 54.684 22.729 51.916 52.340 26.938 50.102 54.315 27.655 51.943
+                 LaMSUM + Proportional Voting
+ Llama 58.392 27.108 55.665 50.649 23.808 48.517 57.137 25.105 54.678 52.129 26.270 49.619 53.800 26.845 50.691
+Mixtral 57.175 27.634 54.600 47.158 23.707 45.989 54.314 26.109 51.072 54.463 30.333 52.058 55.656 27.438 53.134
+GPT 60.440 30.006 58.355 56.296 32.109 55.229 58.708 29.022 56.428 54.855 28.124 52.504 55.921 29.919 53.426
+                 LaMSUM + Borda Count
+ Llama 58.098 26.251 56.092 49.007 23.377 47.612 55.296 24.912 52.637 50.295 26.034 48.265 52.582 25.136 49.978
+Mixtral 56.681 25.156 53.525 46.293 21.928 44.820 52.980 19.953 49.414 53.722 27.817 50.869 53.838 25.767 51.680
+GPT 59.188 27.080 56.585 49.107 24.114 47.250 54.167 22.718 51.847 51.858 27.812 49.761 51.226 24.749 48.956
+Table 9: Table showing metric scores from different LLM models for various datasets.
+The best value per dataset is shown in bold and clearly gpt-4o-mini with proportional voting outperforms all the other methods across all the evaluation measures.
+In this table, Llama, Mixtral and GPT refers to llama-3.1-8B, open-mistral-nemo, and gpt-4o-mini respectively.
+Graphical representation of this table is shown in Figure 4.
+Figure 7: Venn diagram showing the overlap in the gold standard summaries obtained from three annotators (HS1, HS2 and HS3) for 5 cities – (a) City A (b) City B (c) City C (d) City D (e) City E.
+Each annotator selects 50 posts to be included in the summary.
+Note that there can be many posts with similar meaning but different annotators may choose to select different posts for gold summary.
+  City A City B City C City D City E
+Category HS1 HS2 HS3 HS1 HS2 HS3 HS1 HS2 HS3 HS1 HS2 HS3 HS1 HS2 HS3
+ PC1 5 8 4 5 6 6 3 4 6 1 6 1 1 2 2
+PC2 4 4 1 3 2 0 5 8 4 3 5 4 3 3 2
+PC3 4 1 3 8 2 6 5 1 4 2 2 4 10 2 4
+PC4 4 8 9 5 1 5 3 6 9 5 8 5 4 2 10
+PC5 8 8 10 10 10 2 5 19 8 5 8 7 5 10 11
+PC6 11 8 13 7 14 6 6 13 12 14 15 12 13 17 12
+PC7 5 2 7 2 2 1 4 4 6 2 5 1 3 3 5
+PC8 5 3 3 2 4 4 3 3 3 3 7 4 3 5 5
+PC9 14 19 17 10 13 15 14 11 19 12 22 21 12 17 22
+PC10 3 0 1 3 2 1 3 2 2 1 3 4 2 2 3
+PC11 14 7 10 10 16 12 11 16 12 9 8 11 7 11 9
+PC12 3 4 7 6 6 3 3 1 2 6 5 8 3 3 3
+PC13 1 2 1 2 1 1 0 0 0 0 0 0 0 1 0
+PC14 3 0 5 5 9 3 3 3 5 4 6 5 4 1 0
+Table 10: Table demonstrates the count of various harassment cases across gold standard summaries.
+Note that a post can have multiple categories associated with it.
+Refer Table 3 for mapping of the category index (PC) with the category name.
+  Posts
+  One person whom my family rejected for marriage is posting my nude pictures in social media platforms, hacked my email id and forwarding the nude videos and pictures to all the contacts through different different email ids.
+Fake Facebook id using my pics and also fake Instagram accounts using my pics, posting bad things about me and my mother.
+Every day calling and texting me with different numbers.
+Till date he has taken 10 sim cards.
+Torturing my family members every day with 40 different numbers.
+My life has become hell.
+I have lost my job.
+Sole bread winner of the family with 2 aged parents 70+ years.
+Nobody is able to help me.
+She was out shopping at a supermarket when she noticed a 35-40 year old man was taking her pictures/videos.
+Initially, she thought that it might just be a misunderstanding and the man must just be using his phone.
+But later got too suspicious and scary because he was following her where ever she was going.
+She even reported the incident to the staff members of that supermarket but before any actions were taken the man had escaped.
+I was harassed at my workplace in 2015 at ABC Technology Solutions while working as a Programmer Analyst Trainee by a senior employee in the team who attempted to establish physical contact/advances several times and I was unable to react and I later complained to my reporting manager.
+Upon complaining to the HR and my reporting managers they claimed that they know rules pertaining to the sexual harassment act and did not take any corrective/legal action and within a few days, I was asked to give forced resignation/termination.
+This incident took place around 7 30 in the night.
+I took a bus home after my college trip and while I was waiting to collect the ticket from the conductor he touched me in my private part in the upper part of my body.
+I was too confused and scared to speak out something.
+After a while I got a seat in the bus and was continuously yelled at for sitting crossing my legs and was threatened to be thrown out of the bus.
+It was late at night and I just wanted to get home safely
+Table 11: Posts selected by LaMSUM are more detailed and provide a clearer description of the incident.
+Detailed posts enable stakeholders to gain a deeper understanding of the incident’s context, facilitating more informed and effective decision making.